### v2

# Strong solutions for a class of stochastic thermo-magneto-hydrodynamic-type systems with multiplicative noise

Abstract.
We establish the existence and uniqueness of strong solutions, in both the PDE and probabilistic sense, for a broad class of nonlinear stochastic partial differential equations (SPDEs) on a bounded domain $\mathscr{O}\subset\mathbb{R}^{d}$ ($d=2,3$), driven by multiplicative Gaussian noise. The solutions are global in time for $d=2$. This theory simultaneously covers several physically relevant systems, including stochastic convective Brinkman–Forchheimer equations, stochastic magnetohydrodynamics (MHD), stochastic Bénard convection in porous media, stochastic convective dynamo system, stochastic thermo-magneto-micropolar fluids, and stochastic diffusive tropical climate model, for which previous results only provide analytically weak martingale or pathwise solutions. The proof relies on Galerkin approximation and compactness argument. Up to a suitable stopping time, we derive strong moment bounds and verify a Cauchy property for the approximate solutions, in the absence of any inherent cancellation structure. By applying a Gronwall-type lemma for stochastic processes, we establish the existence and uniqueness of maximal strong pathwise solutions, which are global in two spatial dimensions.
1. Introduction
This paper investigates the existence and uniqueness of strong solutions, in the sense of PDEs and probability, for a class of abstract nonlinear stochastic partial differential equations (SPDEs) defined on a bounded domain $\mathscr{O}\subset\mathbb{R}^{d}$, where $d=2,3$. These problems are motivated by thermally-coupled models in magneto-hydrodynamics (MHD), which describe the interplay between thermal, magnetic, and fluid dynamic phenomena under stochastic influences. The problem takes this general form:
${\frac{\mathrm{d}\Phi}{\mathrm{d}t}}+\mathcal{A}\Phi+\mathcal{B}(\Phi,\Phi)+\mathcal{R}(\Phi)$ $=g(\Phi)\dot{W},$ (1.1)
$\Phi(0)$ $=\Phi_{0},$
which describes the time evolution of a state vector $\Phi$. Here, $\mathcal{A}$ is a linear operator representing dissipative effects, $\mathcal{B}$ is a bilinear form capturing convective nonlinearities, and $\mathcal{R}$ is a nonlinear operator accounting for other reactive physical effects. The term $g(\Phi)\dot{W}$ represents a multiplicative Gaussian noise modelling random forcing that depends on the state of the system. To cover a range of physically relevant thermo-magneto-hydrodynamical systems, a set of assumptions on the maps $\mathcal{A}$, $\mathcal{B}$, $\mathcal{R}$, and $g$ will be imposed later. These assumptions ensure that (1.1) encompasses several physically relevant models, such as:
  1. (1)
    the stochastic convective Brinkman–Forchheimer equations [8, 22, 25, 30, 37];
  2. (2)
    the stochastic magnetohydrodynamic (MHD) equations [38, 42, 43, 45, 52];
  3. (3)
    the stochastic Boussinesq model for Bénard convection [11, 14, 46, 52];
  4. (4)
    the stochastic convective dynamo system [9, 10, 47];
  5. (5)
    the stochastic thermo–magneto–micropolar fluid systems [28, 29, 48, 50, 51]; and
  6. (6)
    the stochastic diffusive tropical climate model [19, 33, 53, 54],
among others. These problems will be elaborated further in Section 3.
The works cited above address a broad range of topics, such as the physical derivations of the models, various notions of well-posedness, existence of invariant measures, asymptotic behaviours, and large deviation principles. In what follows, we restrict our attention to the analytical studies concerning the existence of solutions. In particular, for many of the problems listed above, the existence of analytically weak martingale or analytically weak pathwise solutions has been studied in separate works, which we now briefly summarise:
  1. (1)
    Stochastic convective Brinkman–Forchheimer equations: Global martingale weak and pathwise weak solutions are shown to exist in [22] for $d=2,3$ under Dirichlet boundary condition.
  2. (2)
    Stochastic MHD equations: For $d=2,3$, the existence of global martingale weak solutions is established in [43, 52] for bounded domains, and in [38] for more general Poincaré domains. Pathwise uniqueness is also shown for $d=2$, which implies the existence of pathwise weak solutions.
  3. (3)
    Stochastic Boussinesq model for Bénard convection: Local pathwise weak solutions exist in bounded domains for $d=2,3$, as shown in [11], and these solutions are global for $d=2$.
  4. (4)
    Stochastic convective dynamo system: To the best of our knowledge, the existence of solutions for this system has not been established yet.
  5. (5)
    Stochastic thermo–magneto–micropolar fluid: Global martingale weak solutions exist for $d=2,3$, as proven in [50, 51].
  6. (6)
    Stochastic diffusive tropical climate model: Local pathwise weak solutions exist for two-dimensional unbounded domains, as shown in [54].
In contrast, the existence of analytically strong pathwise solutions for these problems in bounded domains with physically relevant boundary conditions is currently unknown. It is worth noting that for the stochastic convective Brinkman–Forchheimer equations, the existence and uniqueness of probabilistically and analytically strong solutions on a torus (or domains with periodic boundary conditions) are established in [30, 37]. For the case of Dirichlet boundary condition, [34, 37] remark that the arguments in [23, 35] contain a gap due to the use of some incorrect identities in several parts of the proof. Moreover, the authors implicitly assume throughout the paper that the Leray projector commutes with the Laplacian, which does not hold under Dirichlet boundary conditions. Consequently, the corresponding existence result for the Dirichlet case is invalid.
The primary focus of this work is the rigorous proof of the existence of analytically strong pathwise solutions for a wide class of thermally-coupled stochastic MHD-type systems for $d=2,3$, with solutions being global in time for $d=2$. To the best of our knowledge, this provides the first strong solution theory for several of these stochastic nonlinear systems, which were previously understood only in the weak sense or not at all. In addition, we develop an abstract framework that unifies the treatment of these models, encompassing both previously studied systems and new physically relevant stochastic nonlinear systems that have not been analysed rigorously before.
The proofs are based on Galerkin approximations combined with compactness arguments, but the analysis requires substantially more delicate estimates than in previous works. In particular, following [24], we derive strong moment bounds and establish a Cauchy property for the approximate solutions up to a suitable stopping time, despite the absence of any inherent cancellation structure in the equations. The final step employs a Gronwall-type lemma for stochastic processes, which allows us to conclude the existence and uniqueness of maximal strong pathwise solutions, with global-in-time existence guaranteed for $d=2$.
This paper is organised as follows:
  • Section 2 sets up the functional framework in which the problem is posed and summarises the standing assumptions used throughout the paper.
  • Section 3 presents several nonlinear physical models that satisfy the assumptions outlined in Section 2.
  • Section 4 establishes the existence and uniqueness of a local strong pathwise solution to problem (1.1).
  • Section 5 extends the local strong solution to a maximal (random) time interval for $d=2,3$, and shows that this maximal strong solution is global when $d=2$.
  • Auxiliary results used in the paper are collected in Appendix A.
Our main results are the existence of a unique maximal strong pathwise solution for $d=2,3$ (Theorem 5.2) and the existence of a global strong pathwise solution for $d=2$ (Theorem 5.3).
2. Preliminaries
2.1. Notations
We begin by defining some notations used in this paper. For $p\in[1,\infty]$ and $k\in\mathbb{N}$, the function spaces $L^{p}$ and $W^{k,p}$ denote, respectively, the Lebesgue space of $p$-th integrable functions and the Sobolev space of $L^{p}$ functions whose weak derivatives up to order $k$ also belong to $L^{p}$, respectively. As usual, we set $H^{k}:=W^{k,2}$.
If $X$ is a Banach space, $L^{p}(0,T;X)$ and $W^{k,p}(0,T;X)$ denote, respectively, the usual Lebesgue and Sobolev spaces of strongly measurable functions on $(0,T)$ taking values in $X$. Similarly, the space $L^{p}(\Omega;X)$ denotes the space of $X$-valued strongly measurable random variables with finite $p$-th moment, where $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. The space $C([0,T];X)$ and $C_{\mathrm{w}}([0,T];X)$ denote the spaces of functions on $[0,T]$ taking values in $X$ that are continuous with respect to the strong and weak topologies of $X$, respectively.
In the analysis, the constant $C$ denotes a generic constant which may take different values at different occurrences. If the dependence of $C$ on a variable, e.g. $s$, is highlighted, we will write $C_{s}$.
2.2. Functional setting
Let $(\mathcal{H},\left|{\cdot}\right|)$ be a separable Hilbert space and let $\mathcal{A}$ be an unbounded, self-adjoint, positive linear operator on $\mathcal{H}$, defined on a domain $\mathrm{D}(\mathcal{A})\subset\mathcal{H}$, such that $\mathcal{A}$ is a bijection onto $\mathcal{H}$ with a compact inverse. Consequently, there exists an orthonormal basis $\{e_{k}\}_{k\in\mathbb{N}}$ for $\mathcal{H}$ consisting of eigenfunctions of $\mathcal{A}$. The associated positive eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{N}}$ form an unbounded increasing sequence. Given $\alpha>0$, let
$\mathrm{D}(\mathcal{A}^{\alpha})=\left\{v\in\mathcal{H}:\sum_{k=1}^{\infty}\lambda_{k}^{2\alpha}\left|{\left\langle v,e_{k}\right\rangle_{\mathcal{H}}}\right|^{2}<\infty\right\}.$
For $v\in\mathrm{D}(\mathcal{A}^{\alpha})$, we define
$\mathcal{A}^{\alpha}v:=\sum_{k=1}^{\infty}\lambda_{k}^{\alpha}\left\langle v,e_{k}\right\rangle e_{k}\quad\text{and}\quad\left|{v}\right|_{\mathrm{D}(\mathcal{A}^{\alpha})}^{2}:=\left|{\mathcal{A}^{\alpha}v}\right|^{2}=\sum_{k=1}^{\infty}\lambda_{k}^{2\alpha}\left|{\left\langle v,e_{k}\right\rangle_{\mathcal{H}}}\right|^{2}.$
Let $\mathcal{V}=\mathrm{D}(\mathcal{A}^{\frac{1}{2}})$, equipped with the norm $\left\|{v}\right\|:=\left|{v}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{2}})}$ for any $v\in\mathcal{V}$. Let $\mathcal{V}^{\prime}$ denote the dual of $\mathcal{V}$ with respect to the inner product $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$ of $\mathcal{H}$. Thus, we have a quartet of compactly embedded Hilbert spaces $\mathrm{D}(\mathcal{A})\subset\mathcal{V}\subset\mathcal{H}\subset\mathcal{V}^{\prime}$. Let $\left\langle u,v\right\rangle$ denote the duality pairing between $u\in\mathcal{V}$ and $v\in\mathcal{V}^{\prime}$, such that $\left\langle u,v\right\rangle=\left\langle u,v\right\rangle_{\mathcal{H}}$ whenever $u\in\mathcal{V}$ and $v\in\mathcal{H}$.
2.3. Assumptions
We assume that the bilinear form $\mathcal{B}$ satisfies the following conditions.
Condition (B):
  • $\mathcal{B}:\mathcal{V}\times\mathcal{V}\to\mathcal{V}^{\prime}$ is a bilinear continuous map.
  • For any $v_{1},v_{2},v_{3}\in\mathcal{V}$,
    $\left\langle\mathcal{B}(v_{1},v_{2}),v_{3}\right\rangle=-\left\langle\mathcal{B}(v_{1},v_{3}),v_{2}\right\rangle.$ (2.1)
  • For any $v_{1},v_{2},v_{3}\in\mathcal{V}$,
    $\left|{\left\langle\mathcal{B}(v_{1},v_{2}),v_{3}\right\rangle}\right|\leq C\left\|{v_{1}}\right\|\left\|{v_{2}}\right\|\left|{v_{3}}\right|^{\frac{1}{2}}\left\|{v_{3}}\right\|^{\frac{1}{2}}.$ (2.2)
  • There exists $C>0$ such that for any $v_{1},v_{3}\in\mathcal{V}$ and $v_{2}\in\mathrm{D}(\mathcal{A})$,
    $\left|{\left\langle\mathcal{B}(v_{1},v_{2}),v_{3}\right\rangle}\right|\leq\begin{cases}C\left|{v_{1}}\right|^{\frac{1}{2}}\left\|{v_{1}}\right\|^{\frac{1}{2}}\left\|{v_{2}}\right\|^{\frac{1}{2}}\left|{\mathcal{A}v_{2}}\right|^{\frac{1}{2}}\left|{v_{3}}\right|,&\text{if $d=2$},\\[4.30554pt] C\left\|{v_{1}}\right\|\left\|{v_{2}}\right\|^{\frac{1}{2}}\left|{\mathcal{A}v_{2}}\right|^{\frac{1}{2}}\left|{v_{3}}\right|,&\text{if $d\leq 3$}.\end{cases}$ (2.3)
  • $\mathcal{B}(v,v)\in\mathrm{D}(\mathcal{A}^{\frac{1}{8}})$ for all $v\in\mathrm{D}(\mathcal{A})$, and there exists $C>0$ such that for any $v\in\mathrm{D}(\mathcal{A})$,
    $\left|{\mathcal{B}(v,v)}\right|^{2}$ $\leq C\left\|{v}\right\|^{3}\left|{\mathcal{A}v}\right|,$ (2.4)
    $\left|{\mathcal{B}(v,v)}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}$ $\leq C\left\|{v}\right\|^{2}\left|{\mathcal{A}v}\right|^{2}.$ (2.5)
In particular, the antisymmetry property (2.1) implies $\left\langle\mathcal{B}(v_{1},v_{2}),v_{2}\right\rangle=0$ for any $v_{1},v_{2}\in\mathcal{V}$. Subsequently, we will write $\mathcal{B}(v):=\mathcal{B}(v,v)$ for simplicity.
We assume that the nonlinear operator $\mathcal{R}$ satisfies the following conditions.
Condition (R):
  • $\mathcal{R}:\mathcal{V}\to\mathcal{H}$ is a continuous map defined by $\mathcal{R}=\mathcal{S}+\mathcal{F}$ where $\mathcal{S},\mathcal{F}:\mathcal{V}\to\mathcal{H}$ satisfy, for any $v\in\mathcal{V}$,
    $\left\langle\mathcal{S}(v),v\right\rangle\geq 0\;\text{ and }\;\left|{\mathcal{F}(v)}\right|\leq C\left(1+\left\|{v}\right\|\right).$ (2.6)
  • There exists $C>0$ such that for any $v\in\mathcal{V}$,
    $\left|{\mathcal{R}(v)}\right|$ $\leq C\left(1+\left\|{v}\right\|^{2}\right)\left\|{v}\right\|.$ (2.7)
    Moreover, for any $v\in\mathrm{D}(\mathcal{A})$,
    $\left|{\mathcal{R}(v)}\right|\leq\begin{cases}C\left(1+\left|{v}\right|^{\frac{3}{2}}\left|{\mathcal{A}v}\right|^{\frac{1}{2}}\right)\left\|{v}\right\|,&\text{if $d=2$},\\[4.30554pt] C\left(1+\left|{v}\right|^{\frac{1}{2}}\left\|{v}\right\|\left|{\mathcal{A}v}\right|^{\frac{1}{2}}\right)\left\|{v}\right\|,&\text{if $d\leq 3$}.\end{cases}$ (2.8)
  • There exists $C>0$ such that for any $v_{1},v_{2}\in\mathcal{V}$
    $\left|{\mathcal{R}(v_{1})-\mathcal{R}(v_{2})}\right|_{\mathcal{H}}\leq C\left(1+\left\|{v_{1}}\right\|^{2}+\left\|{v_{2}}\right\|^{2}\right)\left\|{v_{1}-v_{2}}\right\|$ (2.9)
  • $\mathcal{R}(v)\in\mathrm{D}(\mathcal{A}^{\frac{1}{8}})$ for all $v\in\mathrm{D}(\mathcal{A})$, and there exists $C>0$ such that for any $v\in\mathrm{D}(\mathcal{A})$,
    $\left|{\mathcal{R}(v)}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}$ $\leq C\left(1+\left\|{v}\right\|^{4}\right)\left\|{v}\right\|\left|{\mathcal{A}v}\right|.$ (2.10)
The stochastic term in (1.1) can be written formally as
$g(\Phi)\dot{W}=\sum_{k=1}^{\infty}g_{k}(\Phi)\dot{\beta}_{k},$ (2.11)
where $\beta_{k}$ are independent standard Brownian motions. We need to impose some assumptions on $g=\{g_{k}\}_{k\in\mathbb{N}}$. To this end, we first define the following notion: for any normed spaces $Y$ and $Z$, a progressively measurable map $h:\Omega\times[0,\infty)\times Y\to Z$ is said to be uniformly Lipschitz with constant $L_{Y}$ if, for all $u,v\in Y$ and $(\omega,t)\in\Omega\times[0,\infty)$,
$\left|{h(\omega,t,u)-h(\omega,t,v)}\right|_{Z}\leq L_{Y}\left|{u-v}\right|_{Y}\quad\text{and}\quad\left|{h(\omega,t,u)}\right|_{Z}\leq L_{Y}\left(1+\left|{u}\right|_{Y}\right).$
The collection of all such mappings is denoted by $\text{Lip}(Y,Z)$.
For our analysis, we shall assume the following condition on $g$.
Condition (G): A progressively measurable map $g:\Omega\times[0,\infty)\times\mathcal{H}\to\ell^{2}(\mathcal{H})$ is given such that
$g\in\text{Lip}\big(\mathcal{H},\ell^{2}(\mathcal{H})\big)\cap\text{Lip}\big(\mathcal{V},\ell^{2}(\mathcal{V})\big)\cap\text{Lip}\big(\mathrm{D}(\mathcal{A}),\ell^{2}(\mathrm{D}(\mathcal{A}))\big).$ (2.12)
3. Physical models
We now present several physically relevant models that fall within the scope of our framework. We note, however, that our assumptions are not limited to these thermo-magneto-fluid models and can accommodate a broader class of nonlinear stochastic PDEs arising in various contexts.
In the following sections, $\mathscr{O}\subset\mathbb{R}^{d}$, for $d=2,3$, denotes an open bounded domain with $C^{2}$-smooth boundary or a convex Lipschitz domain.
3.1. Stochastic convective Brinkman–Forchheimer equations
The stochastic convective Brinkman–Forchheimer (or damped Navier–Stokes) equations [8, 25, 30, 37] with Dirichlet (no-slip) boundary conditions read:
$\partial_{t}u-\nu\Delta u+(u\cdot\nabla)u+\alpha u+\beta|u|^{r-1}u+\nabla p=g(u)\dot{W}\;$ in $\mathscr{O}$, (3.1a)
$\mathrm{div}\,u=0\;$ in $\mathscr{O}$, (3.1b)
$u(0,x)=u_{0}(x)\;$ in $\mathscr{O}$, (3.1c)
$u=0\;$ on $\partial\mathscr{O}$, (3.1d)
where $r\in[2,3]$ and $g$ satisfies (2.12). Here, $u:\mathscr{O}\to\mathbb{R}^{d}$ is the fluid velocity and $p:\mathscr{O}\to\mathbb{R}$ is the pressure. The positive coefficients $\nu$, $\alpha$, and $\beta$ respectively denote the kinematic viscosity (Brinkman coefficient), the Darcy coefficient (permeability of porous medium), and the Forchheimer coefficient. The system (3.1) describes the flow of a viscous fluid at sufficiently high velocity and moderate porosity, perturbed by a stochastic forcing term $g(u)\dot{W}$, where $g$ satisfies condition (G) in (2.12).
The existence and uniqueness of probabilistically and analytically strong solutions to (3.1) on a torus (or a domain with periodic boundary conditions) are established in [30, 37]. For the case of Dirichlet boundary condition, a remark is made in [34, 37] that the arguments presented in [23, 35] contain a gap stemming from the use of an incorrect identity in various parts of the proof, which invalidates the corresponding existence result. In this work, we correct these mistakes and establish the existence of pathwise strong solutions for $r\in[2,3]$, including the case of Dirichlet boundary condition. When $d=2$, the solutions are shown to exist globally in time. Our results also extend to the case of Navier-type (or free) boundary conditions [4, 56].
Define the spaces:
$\mathcal{V}$ $:=\{u\in H^{1}_{0}(\mathscr{O}):\mathrm{div}\,u=0\},$
$\mathcal{H}$ $:=\{u\in L^{2}(\mathscr{O}):\mathrm{div}\,u=0\text{ and }\left.u\cdot n\right|_{\partial\mathscr{O}}=0\}.$
In the definition of $\mathcal{H}$ here and in what follows, the condition $\mathrm{div}\,u=0$ is understood in the sense of distributions, while the trace of the normal component of $u$ on the boundary is interpreted in a suitable generalised sense (see [44, Chapter I, Section 1.3] and [20, Section III.2]). By means of the Helmholtz–Leray projector $\mathbb{P}:L^{2}(\mathscr{O})\to\mathcal{H}$, noting that $\mathbb{P}v=v$ and $\mathbb{P}[g(v)\dot{W}]=\sum_{k=1}^{\infty}g_{k}(v)\dot{\beta}_{k}$ due to (2.11) and (2.12), we can project (3.1) onto the space of divergence-free vector fields to formulate problem (3.1) in the form (1.1), with
$\mathcal{A}v=-\nu\mathbb{P}\Delta v,$ $\forall v\in\mathrm{D}(\mathcal{A}),$ (3.2)
$\mathcal{B}(v_{1},v_{2})=\mathbb{P}\big[(v_{1}\cdot\nabla)v_{2}\big],$ $\forall v_{1},v_{2}\in\mathcal{V},$ (3.3)
$\mathcal{R}(v)=\alpha v+\beta\mathbb{P}\big[|v|^{r-1}v\big],$ $\forall v\in\mathcal{V},$ (3.4)
where $\mathrm{D}(\mathcal{A}):=H^{2}(\mathscr{O})\cap\mathcal{V}$. The operator $\mathcal{A}:\mathrm{D}(\mathcal{A})\to\mathcal{H}$ is called the Dirichlet–Stokes operator.
The map $\mathcal{B}$ satisfies (2.1) by the definition of $\mathbb{P}$ and simple calculations. It also satisfies inequalities (2.2) and (2.3) by a standard argument using Hölder’s inequality and Sobolev embedding (see [44]). Furthermore, by the boundedness of $\mathbb{P}$ and Hölder’s inequality (with exponents $p=q=4$ when $d=2$, and $p=6,q=3$ when $d=3$), we have
$\left|{\mathcal{B}(v,v)}\right|^{2}\leq\left|{v}\right|_{L^{p}}^{2}\left|{\nabla v}\right|_{L^{q}}^{2}\leq C\big(\left|{v}\right|\left\|{v}\right\|\big)\big(\left\|{v}\right\|\left|{\mathcal{A}v}\right|\big)\leq C\left\|{v}\right\|^{3}\left|{\mathcal{A}v}\right|,$ (3.5)
where in the penultimate step we used the Gagliardo–Nirenberg inequalities, thus showing (2.4). To show (2.5), we first note that $(v\cdot\nabla)v\in H^{1}(\mathscr{O})$ for any $v\in\mathrm{D}(\mathcal{A})$, since by the Hölder and the Gagliardo–Nirenberg inequalities,
$\left|{(v\cdot\nabla)v}\right|_{H^{1}}^{2}$ $\leq\left|{v}\right|_{L^{4}}^{2}\left|{\nabla v}\right|_{L^{4}}^{2}+\left|{v}\right|_{L^{\infty}}^{2}\left|{\mathcal{A}v}\right|^{2}+\left|{\nabla v}\right|_{L^{p}}^{2}\left|{\nabla v}\right|_{L^{q}}^{2}\leq C\left\|{v}\right\|\left|{\mathcal{A}v}\right|^{3}.$ (3.6)
In particular, this implies $\mathcal{B}(v,v)\in\mathrm{D}(\mathcal{A}^{\frac{1}{8}})$ by the boundedness of $\mathbb{P}$ in $H^{\frac{1}{4}}(\mathscr{O})$ and the characterisation of the domains of fractional powers of the Dirichlet–Stokes operator (see [17, 32]). Furthermore, by interpolation, we infer that
$\left|{\mathcal{B}(v,v)}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}$ $\leq\left|{(v\cdot\nabla)v}\right|_{H^{\frac{1}{2}}}^{2}\leq\left|{(v\cdot\nabla)v}\right|_{L^{2}}\left|{(v\cdot\nabla)v}\right|_{H^{1}}\leq C\left\|{v}\right\|^{2}\left|{\mathcal{A}v}\right|^{2},$ (3.7)
where in the last step we used (3.5) and (3.6). This proves (2.5). Thus, condition (B) in Section 2.3 is satisfied by the map $\mathcal{B}$.
The map $\mathcal{R}$ clearly satisfies (2.6) with $\mathcal{F}(v)=0$. It remains to show that the map satisfies (2.7), (2.8), (2.9), and (2.10). For ease of presentation, we assume $r=3$, noting that similar argument works for $r\leq 3$. First, for any $v\in\mathcal{V}$, we have by Hölder’s inequality and the Sobolev embedding,
$\left|{\mathcal{R}(v)}\right|^{2}\leq C\left(\left|{v}\right|^{2}+\left|{v}\right|_{L^{6}}^{6}\right)\leq C\left(1+\left\|{v}\right\|^{4}\right)\left\|{v}\right\|^{2},$ (3.8)
thus showing (2.7). Next, we show (2.8). By the Hölder inequality (with exponents $p=q=4$ when $d=2$, and $p=6,q=3$ when $d=3$) and the Gagliardo–Nirenberg inequalities, we have for any $v_{1},v_{2}\in\mathcal{V}$,
$\left|{\left\langle\mathcal{R}(v_{1}),v_{2}\right\rangle}\right|$ $\leq C\left(\left|{v_{1}}\right|_{L^{2}}+\left|{v_{1}}\right|_{L^{\infty}}\left|{v_{1}}\right|_{L^{p}}\left|{v_{1}}\right|_{L^{q}}\right)\left|{v_{2}}\right|_{L^{2}}$
$\leq\begin{cases}C\left(\left\|{v_{1}}\right\|+\left|{v_{1}}\right|^{\frac{1}{2}}\left|{\mathcal{A}v_{1}}\right|^{\frac{1}{2}}\left|{v_{1}}\right|\left\|{v_{1}}\right\|\right)\left|{v_{2}}\right|,\;&\text{ if $d=2$,}\\[4.30554pt] C\left(\left\|{v_{1}}\right\|+\left\|{v_{1}}\right\|^{\frac{1}{2}}\left|{\mathcal{A}v_{1}}\right|^{\frac{1}{2}}\left\|{v_{1}}\right\|\left|{v_{1}}\right|^{\frac{1}{2}}\left\|{v_{1}}\right\|^{\frac{1}{2}}\right)\left|{v_{2}}\right|,\;&\text{ if $d=3$},\end{cases}$
as required. By a similar argument, for $d\leq 3$ and any $v_{1},v_{2}\in\mathcal{V}$ and $v_{3}\in\mathcal{H}$, we obtain
$\left|{\mathcal{R}(v_{1})-\mathcal{R}(v_{2})}\right|$ $\leq\left(\left|{v_{1}}\right|_{L^{6}}^{2}+\left|{v_{2}}\right|_{L^{6}}\left|{v_{1}+v_{2}}\right|_{L^{6}}\right)\left|{v_{1}-v_{2}}\right|_{L^{6}}$
$\leq C\left(1+\left\|{v_{1}}\right\|^{2}+\left\|{v_{2}}\right\|^{2}\right)\left\|{v_{1}-v_{2}}\right\|$
by the Sobolev embedding $\mathcal{V}\hookrightarrow L^{6}$ and Young’s inequality. This shows (2.9). Finally, we will show (2.10). To this end, note that $\left|{v}\right|^{2}v\in H^{1}(\mathscr{O})$ for any $v\in\mathrm{D}(\mathcal{A})$ since by Hölder’s inequality and the Sobolev embedding,
$\left|{\left|{v}\right|^{2}v}\right|_{H^{1}}\leq\left|{v}\right|_{L^{6}}^{3}+\left|{v}\right|_{L^{6}}^{2}\left|{\nabla v}\right|_{L^{6}}\leq C\left\|{v}\right\|^{2}\left|{\mathcal{A}v}\right|.$ (3.9)
Hence, by the boundedness of $\mathbb{P}$ in $\mathrm{D}(\mathcal{A}^{\frac{1}{8}})$ and interpolation, noting (3.8) and (3.9), we obtain
$\left|{\mathcal{R}(v)}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}\leq C\left|{v}\right|_{H^{\frac{1}{2}}}^{2}+C\left|{\left|{v}\right|^{2}v}\right|_{H^{\frac{1}{2}}}^{2}$ $\leq C\left|{v}\right|\left\|{v}\right\|+C\left|{\left|{v}\right|^{2}v}\right|_{L^{2}}\left|{\left|{v}\right|^{2}v}\right|_{H^{1}}$
$\leq C\left(1+\left\|{v}\right\|^{4}\right)\left\|{v}\right\|\left|{\mathcal{A}v}\right|,$ (3.10)
proving (2.10). This shows condition (R) in Section 2.3 is satisfied by the map $\mathcal{R}$.
Remark 3.1.
In certain physical applications, it is argued that Navier-type (Hodge or perfect slip) boundary conditions specified by
$u\cdot n=0\quad\text{and}\quad\mathrm{curl}\,u\times n=0\quad\text{on $\partial\mathscr{O}$}$ (3.11)
are more appropriate [1]. In this case, the operator $\mathcal{A}=-\nu\mathbb{P}\Delta$ is the Hodge–Stokes operator, and the corresponding function spaces are
$\mathcal{V}$ $:=\{u\in H^{1}(\mathscr{O}):\mathrm{div}\,u=0\text{ in $\mathscr{O}$, }\left.u\cdot n\right|_{\partial\mathscr{O}}=0\,\text{ and }\left.\mathrm{curl}\,u\times n\right|_{\partial\mathscr{O}}=0\},$
$\mathcal{H}$ $:=\{u\in L^{2}(\mathscr{O}):\mathrm{div}\,u=0\text{ in $\mathscr{O}$ and }\left.u\cdot n\right|_{\partial\mathscr{O}}=0\},$
$\mathrm{D}(\mathcal{A})$ $:=H^{2}(\mathscr{O})\cap\mathcal{V}.$
The verification of condition (B) and condition (R) proceeds in the same manner. To verify (3.7) and (3.1), the corresponding results in [31] for the Hodge–Stokes operator are used.
3.2. Stochastic magnetohydrodynamic equations
The dynamics of turbulent motion of plasmas and liquid metals [6] can be described by the stochastic magnetohydrodynamic (MHD) equations [38, 42], which is a coupled system of stochastic Navier–Stokes and stochastic Maxwell-type equations. Such model is important in the study of the magnetic reconnection problem in astrophysics [39]. The stochastic MHD system read:
$\partial_{t}u-\nu\Delta u+(u\cdot\nabla)u-(B\cdot\nabla)B+\nabla\left(p+\frac{1}{2}\left|{B}\right|^{2}\right)=g_{1}(u,B)\dot{W}_{1}\;$ in $\mathscr{O}$, (3.12a)
$\partial_{t}B-\kappa\Delta B+(u\cdot\nabla)B-(B\cdot\nabla)u=g_{2}(u,B)\dot{W}_{2}\;$ in $\mathscr{O}$, (3.12b)
$\mathrm{div}\,u=\mathrm{div}\,B=0\;$ in $\mathscr{O}$, (3.12c)
$u(0,x)=u_{0}(x),\;B(0,x)=B_{0}(x)\;$ in $\mathscr{O}$, (3.12d)
$u=0,\;$ on $\partial\mathscr{O}$, (3.12e)
$B\cdot n=0,\;\mathrm{curl}\;B\times n=0\;$ on $\partial\mathscr{O}$. (3.12f)
The fields $u:\mathscr{O}\to\mathbb{R}^{d}$ and $B:\mathscr{O}\to\mathbb{R}^{d}$ denote the fluid velocity and the magnetic flux density, respectively. Here, $g_{1}\dot{W}_{1}$ and $g_{2}\dot{W}_{2}$ are the stochastic forcing terms, where $W_{1}$ and $W_{2}$ are independent real-valued Brownian motions. One can also add the term $\alpha u+\beta\left|{u}\right|^{r-1}u$ to the left-hand side of (3.12a) to model a damped flow or a flow in porous media as in [5, 45].
The existence of martingale weak (in the sense of PDEs) solutions to the stochastic MHD systems with multiplicative noise for $d\leq 3$ is shown in [43, 52] for a bounded domain, and in [38] for a Poincaré domain. Pathwise uniqueness of the solution is also derived for $d\leq 2$, which implies the existence of pathwise solutions (weak in the sense of PDEs). The argument in this paper establishes the existence and uniqueness of probabilistically and analytically strong solutions to the stochastic MHD system (including its Brinkman–Forchheimer modification). Our results also extend to the case of perfect slip boundary conditions for the fluid (3.11) or periodic boundary conditions.
The appropriate functional setting for this problem will now be described. Define the spaces:
$V_{1}$ $:=\{u\in H^{1}_{0}(\mathscr{O}):\mathrm{div}\,u=0\},$
$H_{1}$ $:=\{u\in L^{2}(\mathscr{O}):\mathrm{div}\,u=0\text{ in $\mathscr{O}$ and }\left.u\cdot n\right|_{\partial\mathscr{O}}=0\},$
$V_{2}$ $:=\{B\in H^{1}(\mathscr{O}):\mathrm{div}\,B=0\text{ in $\mathscr{O}$, }\left.B\cdot n\right|_{\partial\mathscr{O}}=0\,\text{ and }\left.\mathrm{curl}\,B\times n\right|_{\partial\mathscr{O}}=0\},$
$H_{2}$ $:=H_{1}=\{B\in L^{2}(\mathscr{O}):\mathrm{div}\,u=0\text{ in $\mathscr{O}$ and }\left.u\cdot n\right|_{\partial\mathscr{O}}=0\},$
$\mathcal{V}$ $:=V_{1}\times V_{2},$
$\mathcal{H}$ $:=H_{1}\times H_{2}.$
As before, $\mathbb{P}:L^{2}(\mathscr{O})\to H_{1}$ is the Helmholtz–Leray projector. Let $A_{1}=-\nu\mathbb{P}\Delta$ be the Dirichlet–Stokes operator and $A_{2}=-\kappa\Delta$ be the Hodge Laplacian operator. Let $\mathcal{A}=A_{1}\oplus A_{2}$ with domain $\mathrm{D}(\mathcal{A})=\mathcal{V}\cap\left[H^{2}(\mathscr{O})\right]^{2}$ acting on $\Phi=(u,B)$. The map $\mathcal{B}$ is defined by
$\mathcal{B}(\Phi_{1},\Phi_{2})=\begin{pmatrix}\mathbb{P}\big[(u_{1}\cdot\nabla)u_{2}-(B_{1}\cdot\nabla)B_{2}\big]\\[2.15277pt] (u_{1}\cdot\nabla)B_{2}-(B_{1}\cdot\nabla)u_{2}\end{pmatrix}$
for $\Phi_{i}=(u_{i},B_{i})\in\mathcal{V}$, where $i=1,2$. The stochastic term is $g(\Phi)\dot{W}=(g_{1}\dot{W}_{1},\,g_{2}\dot{W}_{2})^{\top}$. In this case, $\mathcal{R}(\Phi)=0$.
The map $\mathcal{B}$ satisfies (2.1), as well as inequalities (2.2) and (2.3); see [38]. The verification of (2.4) and (2.5) proceeds in the same manner as in (3.5) and (3.7).
3.3. Stochastic Boussinesq model for the Bénard convection
A model for convection in a fluid due to Bénard consists of a coupled system of Navier–Stokes and heat equations, employing Boussinesq approximation (see [2, 18, 21] and references therein). One can also use the Brinkman–Forchheimer equation instead for the fluid flow, as in [3, 41, 46]. Let $e$ be a given vector in $\mathbb{R}^{d}$ and suppose that the non-homogeneous boundary temperature data is fixed. By employing a transformation procedure [26], one can convert the problem into a system with homogeneous Dirichlet data, which reads:
$\partial_{t}u-\nu\Delta u+(u\cdot\nabla)u+\alpha u+\beta|u|^{r-1}u+\nabla p+\theta e=g_{1}(u,\theta)\dot{W}_{1}\;$ in $\mathscr{O}$, (3.13a)
$\partial_{t}\theta-\kappa\Delta\theta+(u\cdot\nabla)\theta+(u\cdot\nabla)\phi=g_{2}(u,\theta)\dot{W}_{2}\;$ in $\mathscr{O}$, (3.13b)
$\mathrm{div}\,u=0\;$ in $\mathscr{O}$, (3.13c)
$u(0,x)=u_{0}(x)\;$ in $\mathscr{O}$, (3.13d)
$u=0,\;\theta=0\;$ on $\partial\mathscr{O}$. (3.13e)
In the above system, $u:\mathscr{O}\to\mathbb{R}^{d}$ and $\theta:\mathscr{O}\to\mathbb{R}$ denote the fluid velocity and the temperature, respectively. The function $\phi\in H^{2}(\mathscr{O})$ is assumed to be given. The non-negative coefficients $\alpha$ and $\beta$ are, respectively, the Darcy and the Forchheimer coefficients. Perfect slip boundary conditions for $u$ as defined by (3.11) can also be used.
The existence of probabilistically strong, analytically weak solutions to the problem (3.13) is established in [11]. The argument in this paper proves the existence and uniqueness of probabilistically and analytically strong solutions to this problem for $d\leq 3$ (which is global when $d\leq 2$).
We now describe the functional setting for this system. Define the spaces:
$V_{1}$ $:=\{v\in H^{1}_{0}(\mathscr{O}):\mathrm{div}\,v=0\},$
$H_{1}$ $:=\{v\in L^{2}(\mathscr{O}):\mathrm{div}\,v=0\;\text{ and }\left.v\cdot n\right|_{\partial\mathscr{O}}=0\},$
$V_{2}$ $:=H^{1}_{0}(\mathscr{O}),$
$H_{2}$ $:=L^{2}(\mathscr{O}),$
$\mathcal{V}$ $:=V_{1}\times V_{2},$
$\mathcal{H}$ $:=H_{1}\times H_{2}.$
Let $A_{1}=-\nu\mathbb{P}\Delta$ be the Dirichlet–Stokes operator and $A_{2}=-\kappa\Delta$ be the Dirichlet Laplacian operator. Let $\mathcal{A}=A_{1}\oplus A_{2}$ with domain $\mathrm{D}(\mathcal{A})=\mathcal{V}\cap\left[H^{2}(\mathscr{O})\right]^{2}$ acting on $\Phi=(u,\theta)$. The map $\mathcal{B}$ is given by
$\mathcal{B}(\Phi_{1},\Phi_{2})=\begin{pmatrix}\mathbb{P}\big[(u_{1}\cdot\nabla)u_{2}\big]\\[2.15277pt] (u_{1}\cdot\nabla)\theta_{2}\end{pmatrix}$
for $\Phi_{i}=(u_{i},\theta_{i})\in\mathcal{V}$, where $i=1,2$. The map $\mathcal{R}$ is defined by
$\mathcal{R}(\Phi)=\begin{pmatrix}\theta e+\alpha u+\mathbb{P}\big[|u|^{r-1}u\big]\\[2.15277pt] (u\cdot\nabla)\phi\end{pmatrix}$,
where $\Phi=(u,\theta)$. The stochastic term is $g(\Phi)\dot{W}=(g_{1}\dot{W}_{1},\,g_{2}\dot{W}_{2})^{\top}$.
By similar argument as in Section 3.1, we infer that $\mathcal{B}$ fulfils condition (B). Next, we verify condition (R) for the map $\mathcal{R}$. Firstly, (2.6) is satisfied with $\mathcal{R}(\Phi)=\mathcal{S}(\Phi)+\mathcal{F}(\Phi)$, where
$\mathcal{S}(\Phi)=\begin{pmatrix}\alpha u+\mathbb{P}\big[|u|^{r-1}u\big]\\[2.15277pt] 0\end{pmatrix},\quad\mathcal{F}(\Phi)=\begin{pmatrix}\theta e\\[2.15277pt] (u\cdot\nabla)\phi\end{pmatrix}$.
Since $\mathcal{S}$ has the same form as (3.4), it remains to verify (2.7), (2.8), (2.9), and (2.10) for the map $\mathcal{F}$. Firstly, by Hölder’s inequality and Sobolev embedding,
$\left|{\mathcal{F}(\Phi)}\right|^{2}$ $\leq C\left|{\theta}\right|^{2}+C\left|{u}\right|_{L^{4}}^{2}\left|{\nabla\phi}\right|_{L^{4}}^{2}\leq C\left\|{\Phi}\right\|^{2},$
which implies (2.7). Next, by Hölder’s inequality and the Sobolev embedding,
$\left|{\left\langle\mathcal{F}(\Phi_{1}),\Phi_{2}\right\rangle}\right|$ $\leq C\left|{\theta_{1}}\right|\left|{u_{2}}\right|+C\left|{u_{1}}\right|_{L^{4}}\left|{\nabla\phi}\right|_{L^{4}}\left|{\theta_{2}}\right|$
$\leq C\left(\left\|{u_{1}}\right\|+\left\|{\theta_{1}}\right\|\right)\left(\left|{u_{2}}\right|+\left|{\theta_{2}}\right|\right)\leq C\left\|{\Phi_{1}}\right\|\left|{\Phi_{2}}\right|,$
which yields (2.8). Inequality (2.9) can be shown similarly. Finally, by interpolation and Gagliardo–Nirenberg inequalities,
$\left|{\mathcal{F}(\Phi)}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}\leq\left|{\mathcal{F}(\Phi)}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{4}})}^{2}$ $\leq C\left|{\mathcal{F}(\Phi)}\right|\left\|{\mathcal{F}(\Phi)}\right\|$
$\leq C\left\|{\Phi}\right\|\big(\left\|{\theta}\right\|+\left|{\nabla u}\right|_{L^{4}}\left|{\nabla\phi}\right|_{L^{4}}+\left|{u}\right|_{L^{\infty}}\left|{\phi}\right|_{H^{2}}\big)$
$\leq C\left\|{\Phi}\right\|\left|{\mathcal{A}\Phi}\right|,$
which implies (2.10), thus completing the verification of condition (R).
3.4. Stochastic convective dynamo system
The dynamo theory [15] is a widely accepted mechanism for the creation and the maintenance of planetary magnetic field. It asserts that the cosmic magnetic field is generated and maintained by the turbulent motion of conducting fluid.
Let $\mathscr{O}=\mathbb{T}^{d-1}\times[0,h]$ be a cylindrical domain, where $h>0$ and $d=2,3$. Let $\{e_{1},\ldots,e_{d}\}$ be the standard basis of $\mathbb{R}^{d}$. We define the following fields:
  1. (i)
    $u:[0,T]\times\mathscr{O}\to\mathbb{R}^{d}$ is the fluid velocity, i.e. $u=(u_{1},\ldots,u_{d})$,
  2. (ii)
    $B:[0,T]\times\mathscr{O}\to\mathbb{R}^{d}$ is the magnetic flux density, i.e. $B=(B_{1},\ldots,B_{d})$,
  3. (iii)
    $\theta:[0,T]\times\mathscr{O}\to\mathbb{R}$ is the temperature profile.
The dynamics of the fluid motion, the magnetic field, and the temperature profile are described by the stochastic convective dynamo system [9, 10, 47]:
$\partial_{t}u-\nu_{1}\Delta u+\sigma e_{d}\times u+(u\cdot\nabla)u-(B\cdot\nabla)B+\nabla\left(p+\frac{1}{2}\left|{B}\right|^{2}\right)+\theta e_{d}=g_{1}\dot{W}_{1}\;$ in $\mathscr{O}$, (3.14a)
$\partial_{t}B-\nu_{2}\Delta B+(u\cdot\nabla)B-(B\cdot\nabla)u=g_{2}\dot{W}_{2}\;$ in $\mathscr{O}$, (3.14b)
$\partial_{t}\theta-\nu_{3}\Delta\theta+(u\cdot\nabla)\theta+e_{d}\cdot u=g_{3}\dot{W}_{3}\;$ in $\mathscr{O}$, (3.14c)
$\mathrm{div}\,u=0,\;\mathrm{div}\,B=0\;$ in $\mathscr{O}$, (3.14d)
$u(0,x)=u_{0}(x),\;B(0,x)=B_{0}(x),\;\theta(0,x)=\theta_{0}(x)\;$ in $\mathscr{O}$, (3.14e)
$u_{d}=B_{d}=0,\;\theta=0\;$ on $\Gamma_{d}$. (3.14f)
The system (3.14) is equipped with a periodic boundary condition in the lateral parts. Here, $\Gamma_{d}:=\{x:x_{d}=0\text{ or }h\}$ is the top and the bottom boundaries. The constant $\sigma$ is the Coriolis force parameter, which is zero when $d=2$.
The functional setting for this system will be described next. Let $H^{1}_{\mathrm{d}}(\mathscr{O})$ denote the closure in $H^{1}(\mathscr{O})$ of the space of compactly supported smooth functions that are periodic in the lateral direction. Define the spaces:
$V_{1}$ $=V_{2}:=\{v\in H^{1}_{\mathrm{d}}(\mathscr{O}):\mathrm{div}\,v=0\},$
$V_{3}$ $:=H^{1}_{\mathrm{d}}(\mathscr{O})$
$H_{1}$ $=H_{2}:=\{v\in L^{2}(\mathscr{O}):\mathrm{div}\,v=0\;\text{ in $\mathscr{O}$ and }\left.v\cdot n\right|_{\Gamma_{d}}=0\},$
$H_{3}$ $:=L^{2}(\mathscr{O}),$
$\mathcal{V}$ $:=V_{1}\times V_{2}\times V_{3},$
$\mathcal{H}$ $:=H_{1}\times H_{2}\times H_{3}.$
For $i=1,2$, let $A_{i}=-\nu_{i}\mathbb{P}\Delta$ be the (periodic-Dirichlet) Stokes operator, and let $A_{3}=-\nu_{3}\Delta$ be the Laplacian operator with the same boundary conditions. Let $\mathcal{A}=A_{1}\oplus A_{2}\oplus A_{3}$ with domain $\mathrm{D}(\mathcal{A})=\mathcal{V}\cap\left[H^{2}(\mathscr{O})\right]^{3}$ acting on $\Phi=(u,B,\theta)$. More generally, it is known that $\mathrm{D}(\mathcal{A}^{\frac{s}{2}})=\big[H^{s}(\mathscr{O})\big]^{3}$ for $s\in[0,\frac{1}{2})$ and $\mathrm{D}(\mathcal{A}^{\frac{1}{2}})=\mathcal{V}$; see [27] for further properties of these function spaces and operators. The map $\mathcal{B}$ is given by
$\mathcal{B}(\Phi_{1},\Phi_{2})=\begin{pmatrix}\mathbb{P}\big[(u_{1}\cdot\nabla)u_{2}-(B_{1}\cdot\nabla)B_{2}\big]\\[2.15277pt] (u_{1}\cdot\nabla)B_{2}-(B_{1}\cdot\nabla)u_{2}\\[2.15277pt] (u_{1}\cdot\nabla)\theta_{2}\end{pmatrix}$
for $\Phi_{i}=(u_{i},B_{i},\theta_{i})\in\mathcal{V}$, where $i=1,2$. The map $\mathcal{R}$ is defined by
$\mathcal{R}(\Phi)=\begin{pmatrix}\sigma\mathbb{P}\big[e_{d}\times u\big]+\theta e_{d}\\[2.15277pt] 0\\[2.15277pt] e_{d}\cdot u\end{pmatrix}$,
where $\Phi=(u,B,\theta)$. The stochastic term is $g(\Phi)\dot{W}=(g_{1}\dot{W}_{1},\,g_{2}\dot{W}_{2},g_{3}\dot{W}_{3})^{\top}$. To show that condition (B) and condition (R) in the assumptions are satisfied, we can proceed in a similar manner as in Section 3.1 and Section 3.3.
3.5. Stochastic micropolar and magneto-micropolar fluid systems
The theory of micropolar fluids [16] describes the behaviour of fluids that exhibit rotational effects and micro-rotational inertia, which include blood and liquid crystal. Taking the electromagnetic effects into consideration, we obtain the stochastic magneto-micropolar fluid system [28, 50, 51], which reads:
$\partial_{t}u-(\mu+\chi)\Delta u+(u\cdot\nabla)u-(B\cdot\nabla)B+\nabla\left(p+\frac{1}{2}\left|{B}\right|^{2}\right)-\chi\,\mathrm{curl}\,w=g_{1}\dot{W}_{1}\;$ in $\mathscr{O}$, (3.15a)
$\partial_{t}w-\gamma\Delta w-(\alpha+\beta)\nabla(\mathrm{div}\,w)+2\chi w+(u\cdot\nabla)w-\chi\,\mathrm{curl}\,u=g_{2}\dot{W}_{2}\;$ in $\mathscr{O}$, (3.15b)
$\partial_{t}B-\nu\Delta B+(u\cdot\nabla)B-(B\cdot\nabla)u=g_{3}\dot{W}_{3}\;$ in $\mathscr{O}$, (3.15c)
$\mathrm{div}\,u=\mathrm{div}\,B=0\;$ in $\mathscr{O}$, (3.15d)
$u(0,x)=u_{0}(x),\;w(0,x)=w_{0}(x),\;B(0,x)=B_{0}(x)\;$ in $\mathscr{O}$, (3.15e)
$u=w=0,\;$ on $\partial\mathscr{O}$, (3.15f)
$B\cdot n=0,\;\mathrm{curl}\;B\times n=0\;$ on $\partial\mathscr{O}$. (3.15g)
Here, $u,w,B$ are vector fields representing the fluid velocity, the micro-rotational vector of the fluid, and the magnetic flux density, respectively. For physical reasons, the coefficients $\mu,\chi,\gamma,\alpha,\beta,\nu$ are assumed to satisfy $\min\{\mu,\chi,\gamma,\nu,\alpha+\beta+\gamma\}>0$. This system can also be further coupled with an equation of the form (3.14c) describing the temperature dynamics [29, 48].
The existence of strong solutions to the deterministic magneto-micropolar fluid system is shown in [40, 49]. In the stochastic case, the existence of martingale weak solutions to the system is proven in [50, 51]. This paper establishes the existence and uniqueness of local pathwise strong solutions to the stochastic magneto-micropolar fluid system for $d\leq 3$. This pathwise strong solution is shown to be global when $d\leq 2$.
The functional setting for this problem will be described next. Define:
$V_{1}$ $:=\{v\in H^{1}_{0}(\mathscr{O}):\mathrm{div}\,v=0\},$
$V_{2}$ $:=H^{1}_{0}(\mathscr{O}),$
$V_{3}$ $:=\{v\in H^{1}(\mathscr{O}):\mathrm{div}\,u=0\text{ in $\mathscr{O}$, }\left.u\cdot n\right|_{\partial\mathscr{O}}=0\,\text{ and }\left.\mathrm{curl}\,u\times n\right|_{\partial\mathscr{O}}=0\},$
$H_{1}$ $:=\{v\in L^{2}(\mathscr{O}):\mathrm{div}\,v=0\;\text{ and }\left.v\cdot n\right|_{\partial\mathscr{O}}=0\},$
$H_{2}$ $:=L^{2}(\mathscr{O})$
$H_{3}$ $:=H_{1}$
$\mathcal{V}$ $:=V_{1}\times V_{2}\times V_{3},$
$\mathcal{H}$ $:=H_{1}\times H_{2}\times H_{3}.$
Let $\mathbb{P}:L^{2}(\mathscr{O})\to H_{1}$ be the Leray–Helmholtz projection, $\Delta_{D}$ be the Dirichlet Laplacian, and $\Delta_{H}$ be the Hodge Laplacian. Define the operators:
$A_{1}$ $:=-(\mu+\chi)\mathbb{P}\Delta_{D},\text{ where }\mathrm{D}(A_{1})=V_{1}\cap H^{2}(\mathscr{O}),$
$A_{2}$ $:=-\gamma\Delta_{D}-(\alpha+\beta)\nabla\mathrm{div}\,,\text{ where }\mathrm{D}(A_{2})=V_{2}\cap H^{2}(\mathscr{O}),$
$A_{3}$ $:=-\nu\Delta_{H},\text{ where }\mathrm{D}(A_{3})=V_{3}\cap H^{2}(\mathscr{O}),$
$\mathcal{A}$ $:=A_{1}\oplus A_{2}\oplus A_{3},\text{ where }\mathrm{D}(\mathcal{A})=A_{1}\times A_{2}\times A_{3}.$
The map $\mathcal{B}$ is given by
$\mathcal{B}(\Phi_{1},\Phi_{2})=\begin{pmatrix}\mathbb{P}\big[(u_{1}\cdot\nabla)u_{2}-(B_{1}\cdot\nabla)B_{2}\big]\\[2.15277pt] (u_{1}\cdot\nabla)w_{2}\\[2.15277pt] (u_{1}\cdot\nabla)B_{2}-(B_{1}\cdot\nabla)u_{2}\end{pmatrix}$
for $\Phi_{i}=(u_{i},w_{i},B_{i})\in\mathcal{V}$, where $i=1,2$. The map $\mathcal{R}$ is defined by
$\mathcal{R}(\Phi)=\begin{pmatrix}-\chi\,\mathrm{curl}\,w\\[2.15277pt] 2\chi w-\chi\,\mathrm{curl}\,u\\[2.15277pt] 0\end{pmatrix}$,
where $\Phi=(u,w,B)$. The stochastic term is $g(\Phi)\dot{W}=(g_{1}\dot{W}_{1},\,g_{2}\dot{W}_{2},g_{3}\dot{W}_{3})^{\top}$. As before, we can verify that our standing assumptions are valid for the operators defined above.
3.6. Stochastic diffusive tropical climate model
A tropical climate model of the equatorial atmosphere is proposed by Frierson–Majda–Pauluis in [19] by performing a Galerkin truncation to the hydrostatic Boussinesq equation. Since then, several physically reasonable modifications of the model have been proposed, including adding diffusion or damping terms to some or all of the equations [33, 53]. Here, we consider the stochastic diffusive tropical climate model:
$\partial_{t}u-\nu_{1}\Delta u+(u\cdot\nabla)u+\nabla p+\mathrm{div}\,(v\otimes v)=g_{1}\dot{W}_{1}\;$ in $\mathscr{O}$, (3.16a)
$\partial_{t}v-\nu_{2}\Delta v+(u\cdot\nabla)v+\nabla\theta+(v\cdot\nabla)u=g_{2}\dot{W}_{2}\;$ in $\mathscr{O}$, (3.16b)
$\partial_{t}\theta-\nu_{3}\Delta\theta+(u\cdot\nabla)\theta+\mathrm{div}\,v=g_{3}\dot{W}_{3}\;$ in $\mathscr{O}$, (3.16c)
$\mathrm{div}\,u=0\;$ in $\mathscr{O}$, (3.16d)
$u(0,x)=u_{0}(x),\;v(0,x)=v_{0}(x),\;\theta(0,x)=\theta_{0}(x)\;$ in $\mathscr{O}$, (3.16e)
$u=v=0,\;\theta=0,\;$ on $\partial\mathscr{O}$, (3.16f)
where $\nu_{1},\nu_{2},\nu_{3}>0$ are diffusivity constants, and
  1. (i)
    $u:[0,T]\times\mathscr{O}\to\mathbb{R}^{d}$ is the barotropic mode,
  2. (ii)
    $v:[0,T]\times\mathscr{O}\to\mathbb{R}^{d}$ is the first baroclinic mode of the velocity vector,
  3. (iii)
    $\theta:[0,T]\times\mathscr{O}\to\mathbb{R}$ is the temperature profile.
In the deterministic case, the system (3.16) is widely studied. Global existence of strong solutions for $d=2$ is established in [33], even assuming $\nu_{3}=0$. Global existence of solutions for $d=2$ with $\nu_{1}\Delta u$ replaced by the fractional Laplacian is shown in [53]; see also [12, 13]. For $d=3$, the system with additional damping terms is studied in [55]. In the stochastic case, local existence and uniqueness of strong solutions for the problem when $d=2$ is shown in [54]. In this paper, we establish the local existence and uniqueness of pathwise strong solutions for $d\leq 3$, which is global when $d=2$.
We describe the functional setting for this problem as follows. Define:
$V_{1}$ $:=\{v\in H^{1}_{0}(\mathscr{O}):\mathrm{div}\,v=0\},$
$V_{2}$ $=V_{3}:=H^{1}_{0}(\mathscr{O}),$
$H_{1}$ $:=\{v\in L^{2}(\mathscr{O}):\mathrm{div}\,v=0\;\text{ and }\left.v\cdot n\right|_{\partial\mathscr{O}}=0\},$
$H_{2}$ $=H_{3}:=L^{2}(\mathscr{O})$
$\mathcal{V}$ $:=V_{1}\times V_{2}\times V_{3},$
$\mathcal{H}$ $:=H_{1}\times H_{2}\times H_{3}.$
Let $A_{1}=-\nu_{1}\mathbb{P}\Delta$ be the Dirichlet–Stokes operator, $A_{2}=-\nu_{2}\Delta$, and $A_{3}=-\nu_{3}\Delta$, where $\Delta$ is Dirichlet Laplacian operator. Let $\mathcal{A}=A_{1}\oplus A_{2}\oplus A_{3}$ with domain $\mathrm{D}(\mathcal{A})=\mathcal{V}\cap\left[H^{2}(\mathscr{O})\right]^{3}$ acting on $\Phi=(u,v,\theta)$. The map $\mathcal{B}$ is given by
$\mathcal{B}(\Phi_{1},\Phi_{2})=\begin{pmatrix}\mathbb{P}\big[(u_{1}\cdot\nabla)u_{2}+\mathrm{div}\,(v_{1}\otimes v_{2})\big]\\[2.15277pt] (u_{1}\cdot\nabla)v_{2}+(v_{1}\cdot\nabla)u_{2}\\[2.15277pt] (u_{1}\cdot\nabla)\theta_{2}\end{pmatrix}$
for $\Phi_{i}=(u_{i},v_{i},\theta_{i})\in\mathcal{V}$, where $i=1,2$. The map $\mathcal{R}$ is defined by
$\mathcal{R}(\Phi)=\begin{pmatrix}0\\[2.15277pt] \nabla\theta\\[2.15277pt] \mathrm{div}\,v\end{pmatrix}$,
where $\Phi=(u,w,B)$. The stochastic term is $g(\Phi)\dot{W}=(g_{1}\dot{W}_{1},\,g_{2}\dot{W}_{2},g_{3}\dot{W}_{3})^{\top}$. We can verify that our standing assumptions are valid for the operators defined above.
4. Existence and uniqueness of a local strong solution
Suppose that we fix a stochastic basis $\mathcal{S}:=(\Omega,\mathcal{F},\mathbb{F},\mathbb{P},\{\beta_{k}\}_{k\in\mathbb{N}})$, where $\mathbb{F}=\{\mathcal{F}_{t}\}_{t\geq 0}$ is a filtration satisfying the usual conditions and $\{\beta_{k}\}_{k\in\mathbb{N}}$ is a sequence of independent Brownian motions adapted to this filtration. We have the following definitions of weak and strong pathwise solutions of (1.1).
Definition 4.1 (local weak pathwise solution).
The pair $(\Phi,\tau)$ is a local weak pathwise solution of (1.1) with initial data $\Phi_{0}$ if $\tau$ is a strictly positive stopping time and $\Phi(\,\cdot\wedge\tau)$ is a predictable process in $\mathcal{V}^{\prime}$ such that
$\Phi(\,\cdot\wedge\tau)\in L^{2}\big(\Omega;C_{\mathrm{w}}([0,\infty);\mathcal{H})\big),\quad\Phi\mathds{1}_{t\leq\tau}\in L^{2}\big(\Omega;L^{2}_{\mathrm{loc}}([0,\infty);\mathcal{V})\big),$ (4.1)
and that for any $t>0$,
$\Phi(t\wedge\tau)+\int_{0}^{t\wedge\tau}\big(\mathcal{A}\Phi+\mathcal{B}(\Phi)+\mathcal{R}(\Phi)\big)\,\mathrm{d}s=\Phi_{0}+\int_{0}^{t\wedge\tau}g(\Phi)\,\mathrm{d}W\quad\text{in $\mathcal{V}^{\prime}$}.$ (4.2)
Definition 4.2 (local strong pathwise solution).
A local weak pathwise solution $(\Phi,\tau)$ is said to be a local strong pathwise solution of (1.1) with initial data $\Phi_{0}$ if $\Phi(\,\cdot\wedge\tau)$ is a predictable process in $\mathcal{H}$ such that
$\Phi(\,\cdot\wedge\tau)\in L^{2}\big(\Omega;C([0,\infty);\mathcal{V})\big),\quad\Phi\mathds{1}_{t\leq\tau}\in L^{2}\big(\Omega;L^{2}_{\mathrm{loc}}([0,\infty);\mathrm{D}(\mathcal{A}))\big),$ (4.3)
and that for any $t>0$, the equality (4.2) holds in $\mathcal{H}$.
Definition 4.3 (maximal strong pathwise solution).
The pair $(\Phi,\xi)$ is said to be a maximal strong pathwise solution if there exists a strictly increasing sequence of stopping times $\{\tau_{n}\}_{n\in\mathbb{N}}$ with $\tau_{n}\uparrow\xi$ a.s., such that each pair $(\Phi,\tau_{n})$ is a local pathwise strong solution and
$\sup_{t<\xi}\left\|{\Phi(t)}\right\|^{2}+\int_{0}^{\xi}\left|{\mathcal{A}\Phi(s)}\right|^{2}\mathrm{d}s=\infty\;\text{ a.s. on the set }\{\xi<\infty\}.$
The sequence $\{\tau_{n}\}_{n\in\mathbb{N}}$ is said to announce (finite time blow up at) $\xi$.
Such a maximal strong pathwise solution $(\Phi,\xi)$ is global if $\mathbb{P}\left[\xi=\infty\right]=1$.
Remark 4.4.
Clearly, the following holds:
  1. (1)
    $\xi$ is a stopping time;
  2. (2)
    we can choose the announcing sequence $\{\tau_{n}\}$ such that
$\sup_{t\leq\tau_{n}}\left\|{\Phi(t)}\right\|^{2}+\int_{0}^{\tau_{n}}\left|{\mathcal{A}\Phi(s)}\right|^{2}\mathrm{d}s=n\;\text{ a.s. on the set }\{\xi<\infty\}.$
To show the existence of a solution to (1.1), we use the Galerkin method and a compactness argument utilising Theorem A.1. Let $\mathcal{H}_{n}:=\text{span}\{e_{1},e_{2},\ldots,e_{n}\}$, where $\{e_{k}\}_{k\in\mathbb{N}}$ is an orthonormal basis for $\mathcal{H}$ consisting of eigenfunctions of $\mathcal{A}$. Let $P_{n}:\mathcal{H}\to\mathcal{H}_{n}$ be the projection operator and let $Q_{n}=I-P_{n}$, where $I$ is the identity operator.
The Galerkin approximation of order $n$ for (1.1) can be written as a problem in $\mathcal{H}_{n}$: find an adapted process $\Phi^{n}\in C([0,T];\mathcal{H}_{n})$ satisfying
$\mathrm{d}\Phi^{n}+\big(\mathcal{A}\Phi^{n}+P_{n}\mathcal{B}(\Phi^{n})+P_{n}\mathcal{R}(\Phi^{n})\big)\mathrm{d}t$ $=\sum_{k=1}^{\infty}P_{n}g_{k}(\Phi^{n})\,\mathrm{d}\beta_{k},$ (4.4)
$\Phi^{n}(0)$ $=\Phi_{0}^{n}=P_{n}\Phi_{0}.$
The existence of a unique solution to (4.4) follows by a standard argument. An important ingredient needed to apply Theorem A.1 is conferred by the following proposition.
Proposition 4.5.
Let $\{\Phi^{n}\}_{n\in\mathbb{N}}$ be a sequence of solutions to (4.4). Suppose that $\Phi_{0}\in\mathcal{V}$. For any $M>1$ and $T>0$, consider a collection of stopping times
$\mathcal{T}_{n}^{M,T}=\left\{\tau\leq T:\left(\sup_{t\in[0,\tau]}\left\|{\Phi^{n}(t)}\right\|^{2}+\int_{0}^{\tau}\left|{\mathcal{A}\Phi^{n}(s)}\right|^{2}\mathrm{d}s\right)^{\frac{1}{2}}\leq M+\left\|{\Phi_{0}^{n}}\right\|\right\},$ (4.5)
and let $\mathcal{T}_{n,m}^{M,T}:=\mathcal{T}_{n}^{M,T}\cap\mathcal{T}_{m}^{M,T}$. Furthermore, for any $n\in\mathbb{N},S>0$, and stopping time $\tau$, let
$A_{n}(\tau,S):=\left\{\sup_{t\in[0,\tau\wedge S]}\left\|{\Phi^{n}(t)}\right\|^{2}+\int_{0}^{\tau\wedge S}\left|{\mathcal{A}\Phi^{n}(s)}\right|^{2}\mathrm{d}s>\left\|{\Phi^{n}_{0}}\right\|^{2}+(M-1)^{2}\right\}.$ (4.6)
Then the following statements hold true:
  1. (i)
    For any $T>0$ and $M>1$,
    $\lim_{n\to\infty}\sup_{m>n}\,\sup_{\tau\in\mathcal{T}_{m,n}^{M,T}}\mathbb{E}\left[\sup_{t\in[0,\tau]}\left\|{\Phi^{m}(t)-\Phi^{n}(t)}\right\|^{2}+\int_{0}^{\tau}\left|{\mathcal{A}\big(\Phi^{m}(s)-\Phi^{n}(s)\big)}\right|^{2}\mathrm{d}s\right]=0.$ (4.7)
  2. (ii)
    For any $T>0$ and $M>1$,
    $\lim_{S\to 0}\sup_{n\in\mathbb{N}}\,\sup_{\tau\in\mathcal{T}_{n}^{M,T}}\mathbb{P}\left[A_{n}(\tau,S)\right]=0.$ (4.8)
Proof.
For $m>n$, let $\Phi^{m,n}:=\Phi^{m}-\Phi^{n}$. We subtract equation (4.4) for $\Phi^{n}$ from that for $\Phi^{m}$ to obtain
$\mathrm{d}\Phi^{m,n}+\mathcal{A}\Phi^{m,n}\mathrm{d}t$ $=\left[P_{n}\mathcal{B}(\Phi^{n})-P_{m}\mathcal{B}(\Phi^{m})+P_{n}\mathcal{R}(\Phi^{n})-P_{m}\mathcal{R}(\Phi^{m})\right]\mathrm{d}t$
$\quad+\sum_{k=1}^{\infty}\left[P_{m}g_{k}(\Phi^{m})-P_{n}g_{k}(\Phi^{n})\right]\mathrm{d}\beta_{k},$
$\Phi^{m,n}(0)$ $=(P_{m}-P_{n})\Phi_{0}.$
By Itô’s lemma applied to $\left\|{\Phi^{m,n}}\right\|^{2}$, we have
$\mathrm{d}\left\|{\Phi^{m,n}}\right\|^{2}+2\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}\mathrm{d}t$
$=2\left\langle P_{n}\mathcal{B}(\Phi^{n})-P_{m}\mathcal{B}(\Phi^{m}),\mathcal{A}\Phi^{m,n}\right\rangle\mathrm{d}t+2\left\langle P_{n}\mathcal{R}(\Phi^{n})-P_{m}\mathcal{R}(\Phi^{m}),\mathcal{A}\Phi^{m,n}\right\rangle\mathrm{d}t$
$\quad+\sum_{k=1}^{\infty}\left\|{P_{m}g_{k}(\Phi^{m})-P_{n}g_{k}(\Phi^{n})}\right\|^{2}\mathrm{d}t+2\sum_{k=1}^{\infty}\left\langle P_{m}g_{k}(\Phi^{m})-P_{n}g_{k}(\Phi^{n}),\mathcal{A}\Phi^{m,n}\right\rangle\mathrm{d}\beta_{k}.$ (4.9)
Let $\tau\in\mathcal{T}_{m,n}^{M,T}$ and let $\tau_{a}$ and $\tau_{b}$ be stopping times such that $0\leq\tau_{a}\leq\tau_{b}\leq\tau$. Integrating (4) over $\tau_{a},r]$, then taking supremum over $r\in[\tau_{a},\tau_{b}]$ and applying the expected value, we obtain by the triangle inequality,
$\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi^{m,n}(t)}\right\|^{2}\right]+2\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}(t)}\right|^{2}\mathrm{d}t\right]$
$\leq\mathbb{E}\left[\left\|{\Phi^{m,n}(\tau_{a})}\right\|^{2}\right]+2\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\left\langle P_{m}\mathcal{B}(\Phi^{m})-P_{n}\mathcal{B}(\Phi^{n}),\mathcal{A}\Phi^{m,n}\right\rangle}\right|\mathrm{d}t\right]$
$\quad+\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\left\langle P_{m}\mathcal{R}(\Phi^{m})-P_{n}\mathcal{R}(\Phi^{n}),\mathcal{A}\Phi^{m,n}\right\rangle}\right|\mathrm{d}t\right]+\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\sum_{k=1}^{\infty}\left\|{P_{m}g_{k}(\Phi^{m})-P_{n}g_{k}(\Phi^{n})}\right\|^{2}\mathrm{d}t\right]$
$\quad+2\mathbb{E}\left[\sup_{r\in[\tau_{a},\tau_{b}]}\left|\sum_{k=1}^{\infty}\int_{\tau_{a}}^{r}\left\langle P_{m}g_{k}(\Phi^{m})-P_{n}g_{k}(\Phi^{n}),\mathcal{A}\Phi^{m,n}\right\rangle\,\mathrm{d}\beta_{k}\right|\right]$
$=:\mathbb{E}\left[\left\|{\Phi^{m,n}(\tau_{a})}\right\|^{2}\right]+I_{1}+\ldots+I_{4}.$ (4.10)
Each term $I_{j}$, $j=1,2,3,4$, will be estimated next. Firstly, we consider the term $I_{1}$. Note that we can write
$P_{m}\mathcal{B}(\Phi^{m})-P_{n}\mathcal{B}(\Phi^{n})=P_{m}\mathcal{B}(\Phi^{m,n},\Phi^{m})+P_{m}\mathcal{B}(\Phi^{n},\Phi^{m,n})+(P_{m}-P_{n})\mathcal{B}(\Phi^{n}).$
By (2.3), we have
$\left|{\left\langle\mathcal{B}(\Phi^{m,n},\Phi^{m}),\mathcal{A}\Phi^{m,n}\right\rangle}\right|$ $\leq C\left\|{\Phi^{m,n}}\right\|\left|{\mathcal{A}\Phi^{m}}\right|\left|{\mathcal{A}\Phi^{m,n}}\right|$
$\leq C\left\|{\Phi^{m,n}}\right\|^{2}\left|{\mathcal{A}\Phi^{m}}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2},$ (4.11)
where in the last step we used Young’s inequality. Similarly, we also have
$\left|{\left\langle\mathcal{B}(\Phi^{n},\Phi^{m,n}),\mathcal{A}\Phi^{m,n}\right\rangle}\right|$ $\leq C\left\|{\Phi^{n}}\right\|\left\|{\Phi^{m,n}}\right\|^{\frac{1}{2}}\left|{\mathcal{A}\Phi^{m,n}}\right|^{\frac{3}{2}}$
$\leq C\left\|{\Phi^{m,n}}\right\|^{2}\left\|{\Phi^{n}}\right\|^{4}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2},$ (4.12)
Furthermore, by Young’s inequality, Lemma A.3, and estimate (2.5), we obtain
$\left|{\left\langle(P_{m}-P_{n})\mathcal{B}(\Phi^{n}),\mathcal{A}\Phi^{m,n}\right\rangle}\right|$ $\leq C\left|{Q_{n}\mathcal{B}(\Phi^{n})}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-\frac{1}{4}}\left|{Q_{n}\mathcal{B}(\Phi^{n})}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-\frac{1}{4}}\left\|{\Phi^{n}}\right\|^{2}\left|{\mathcal{A}\Phi^{n}}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}.$ (4.13)
Combining (4), (4), and (4), we deduce that for the term $I_{1}$ in (4),
$\left|{I_{1}}\right|$ $\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\Big(\left|{\mathcal{A}\Phi^{m}}\right|^{2}+\left\|{\Phi^{n}}\right\|^{4}\Big)+\lambda_{n}^{-\frac{1}{2}}\left\|{\Phi^{n}}\right\|^{2}\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t\right]+\frac{1}{3}\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}\mathrm{d}t\right].$ (4.14)
Next, we consider the term $I_{2}$ in (4). To this end, we write
$P_{m}\mathcal{R}(\Phi^{m})-P_{n}\mathcal{R}(\Phi^{n})=P_{m}\left[\mathcal{R}(\Phi^{m})-\mathcal{R}(\Phi^{n})\right]+(P_{m}-P_{n})\mathcal{R}(\Phi^{n}).$
By (2.8) for $d\leq 3$ and Young’s inequality, we have
$\left|{\left\langle\mathcal{R}(\Phi^{m})-\mathcal{R}(\Phi^{n}),\mathcal{A}\Phi^{m,n}\right\rangle}\right|$ $\leq C\left(1+\left\|{\Phi^{m}}\right\|^{4}+\left\|{\Phi^{n}}\right\|^{4}\right)\left\|{\Phi^{m,n}}\right\|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}.$ (4.15)
Moreover, by an argument similar to that in (4), we obtain
$\left|{\left\langle(P_{m}-P_{n})\mathcal{R}(\Phi^{n}),\mathcal{A}\Phi^{m,n}\right\rangle}\right|$ $\leq C\left|{Q_{n}\mathcal{R}(\Phi^{n})}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-\frac{1}{4}}\left|{Q_{n}\mathcal{R}(\Phi^{n})}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-\frac{1}{4}}\left(1+\left\|{\Phi^{n}}\right\|^{4}\right)\left|{\mathcal{A}\Phi^{n}}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}.$ (4.16)
Thus, we deduce an estimate for the term $I_{2}$ in (4):
$\left|{I_{2}}\right|$ $\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\left(1+\left\|{\Phi^{m}}\right\|^{4}+\left\|{\Phi^{n}}\right\|^{4}\right)\mathrm{d}t\right]$
$\quad+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-\frac{1}{4}}\left(1+\left\|{\Phi^{n}}\right\|^{4}\right)\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t\right]+\frac{1}{3}\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}\mathrm{d}t\right].$ (4.17)
The term $I_{3}$ in (4) will be estimated next. To this end, note that by the triangle inequality, the Lipschitz property of $g$ assumed in Condition (G), and Lemma A.3, we have
$\sum_{k=1}^{\infty}\left\|{P_{m}g_{k}(\Phi^{m})-P_{n}g_{k}(\Phi^{n})}\right\|^{2}$ $\leq\sum_{k=1}^{\infty}\left\|{g_{k}(\Phi^{m})-g_{k}(\Phi^{n})}\right\|^{2}+\left\|{Q_{n}g_{k}(\Phi^{n})}\right\|^{2}$
$\leq C\left\|{\Phi^{m,n}}\right\|^{2}+C\lambda_{n}^{-1}\sum_{k=1}^{\infty}\left|{\mathcal{A}g_{k}(\Phi^{n})}\right|^{2}$
$\leq C\left\|{\Phi^{m,n}}\right\|^{2}+C\lambda_{n}^{-1}\big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\big).$ (4.18)
Therefore,
$\left|{I_{3}}\right|$ $\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left(\left\|{\Phi^{m,n}}\right\|^{2}+\lambda_{n}^{-1}\big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\big)\right)\mathrm{d}t\right].$ (4.19)
The term $I_{4}$ is a stochastic term, which we will estimate using the Burkholder–Davis–Gundy inequality and (4) to obtain
$\left|{I_{4}}\right|$ $\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\sum_{k=1}^{\infty}\left|{\left\langle P_{m}g_{k}(\Phi^{m})-P_{n}g_{k}(\Phi^{n}),\mathcal{A}\Phi^{m,n}\right\rangle}\right|^{2}\mathrm{d}t\right]^{\frac{1}{2}}$
$\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\sum_{k=1}^{\infty}\left\|{P_{m}g_{k}(\Phi^{m})-P_{n}g_{k}(\Phi^{n})}\right\|^{2}\mathrm{d}t\right]^{\frac{1}{2}}$
$\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\left(\left\|{\Phi^{m,n}}\right\|^{2}+\lambda_{n}^{-1}\big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\big)\right)\mathrm{d}t\right]^{\frac{1}{2}}$
$\leq\frac{1}{2}\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi^{m,n}}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left(\left\|{\Phi^{m,n}}\right\|^{2}+\lambda_{n}^{-1}\big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\big)\right)\mathrm{d}t\right],$ (4.20)
where in the last step we applied Young’s inequality. Altogether, from estimates (4.14), (4), (4.19), and (4), we continue from (4) to infer that
$\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi^{m,n}(t)}\right\|^{2}\right]+\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}(t)}\right|^{2}\mathrm{d}t\right]$
$\leq C\mathbb{E}\left[\left\|{\Phi^{m,n}(\tau_{a})}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-\frac{1}{4}}\Big(1+\left\|{\Phi^{n}}\right\|^{4}\Big)\Big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right]$
$\quad+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\Big(1+\left\|{\Phi^{m}}\right\|^{4}+\left\|{\Phi^{n}}\right\|^{4}+\left|{\mathcal{A}\Phi^{m}}\right|^{2}+\lambda_{n}^{-\frac{1}{4}}\left\|{\Phi^{n}}\right\|^{2}\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right]$
$\leq C\mathbb{E}\left[\left\|{\Phi^{m,n}(\tau_{a})}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\Big(1+\left|{\mathcal{A}\Phi^{m}}\right|^{2}\Big)\mathrm{d}t\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-\frac{1}{4}}\Big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right],$
where in the last step we used the fact that $\tau_{a},\tau_{b}\in\mathcal{T}_{m}^{M,T}\cap\mathcal{T}_{n}^{M,T}$ and (4.5) to bound $\left\|{\Phi^{m}}\right\|$ and $\left\|{\Phi^{n}}\right\|$. Here, $C$ depends on $M$, but is independent of $m$, $n$, $\tau_{a}$, or $\tau_{b}$. By the same token, we have
$\int_{0}^{\tau}\left(1+\left|{\mathcal{A}\Phi^{m}(t)}\right|^{2}\right)\mathrm{d}t\leq\left(M+\left\|{\Phi^{0}}\right\|\right)^{2}+T,\;\text{ a.s.}$
We are now in the position to apply the Gronwall lemma for stochastic processes (Lemma A.4) with $b=1+\left|{\mathcal{A}\Phi^{m}}\right|^{2}$, $u=\left\|{\Phi^{m,n}}\right\|^{2}$, $v=\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$, and $w=\lambda_{n}^{-\frac{1}{4}}\big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\big)$. This results in
$\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi^{m,n}}\right\|^{2}\right]+\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}\mathrm{d}t\right]\leq C\mathbb{E}\left[\left\|{\Phi^{0}}\right\|^{2}\right]+C\lambda_{n}^{-\frac{1}{4}},$
where $C$ depends on $M$ and $T$, but is independent of $m$, $n$, or $\tau$. Thus, taking supremum over $\tau\in\mathcal{T}_{m,n}^{M,T}$, then supremum over all $m>n$, and letting $n\to\infty$, we obtain (4.7).
Next, we aim to show (4.8). To this end, by Itô’s lemma applied to $\left\|{\Phi^{n}}\right\|$, we have
$\mathrm{d}\left\|{\Phi^{n}}\right\|^{2}+2\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t$ $=\Big(-2\left\langle\mathcal{B}(\Phi^{n}),\mathcal{A}\Phi^{n}\right\rangle+2\left\langle\mathcal{R}(\Phi^{n}),\mathcal{A}\Phi^{n}\right\rangle+\sum_{k=1}^{\infty}\left\|{P_{n}g_{k}(\Phi^{n})}\right\|^{2}\Big)\mathrm{d}t$
$\quad+2\sum_{k=1}^{\infty}\left\langle g_{k}(\Phi^{n}),\mathcal{A}\Phi^{n}\right\rangle\mathrm{d}\beta_{k}.$ (4.21)
Let $\tau\in\mathcal{T}_{n}^{M,T}$ and $S>0$. Integrating (4) over $[0,t]$, then taking supremum over $t\in[0,S\wedge\tau]$ and applying the expected value, we obtain
$\mathbb{E}\left[\sup_{t\in[0,S\wedge\tau]}\left\|{\Phi^{n}}\right\|^{2}\right]+2\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t\right]$
$\leq\left\|{\Phi_{0}}\right\|^{2}-2\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left|{\left\langle\mathcal{B}(\Phi^{n}),\mathcal{A}\Phi^{n}\right\rangle}\right|\mathrm{d}t\right]+2\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left|{\left\langle\mathcal{R}(\Phi^{n}),\mathcal{A}\Phi^{n}\right\rangle}\right|\mathrm{d}t\right]$
$\quad+\mathbb{E}\left[\int_{0}^{S\wedge\tau}\sum_{k=1}^{\infty}\left\|{P_{m}g_{k}(\Phi^{n})}\right\|^{2}\mathrm{d}t\right]+2\mathbb{E}\left[\sup_{r\in[0,S\wedge\tau]}\left|{\sum_{k=1}^{\infty}\int_{0}^{r}2\left\langle g_{k}(\Phi^{n}),\Phi^{n}\right\rangle\mathrm{d}\beta_{k}}\right|\right].$
$=:\left\|{\Phi_{0}}\right\|^{2}+J_{1}+J_{2}+J_{3}+J_{4}.$ (4.22)
We need to estimate each term on the last line. For the term $J_{1}$, by (2.3) for $d\leq 3$ and Young’s inequality, we have
$\left|{J_{1}}\right|$ $\leq C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left\|{\Phi^{n}}\right\|^{\frac{3}{2}}\left|{\mathcal{A}\Phi^{n}}\right|^{\frac{3}{2}}\mathrm{d}t\right]\leq C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left\|{\Phi^{n}}\right\|^{6}\mathrm{d}t\right]+\frac{1}{4}\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t\right].$ (4.23)
Next, for the term $J_{2}$, applying (2.8) and Young’s inequality yields
$J_{2}$ $\leq C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left(1+\left|{\Phi^{n}}\right|^{\frac{1}{2}}\left\|{\Phi^{n}}\right\|\left|{\mathcal{A}\Phi^{n}}\right|^{\frac{1}{2}}\right)\left\|{\Phi^{n}}\right\|\left|{\mathcal{A}\Phi^{n}}\right|\mathrm{d}t\right]$
$\leq C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left\|{\Phi^{n}}\right\|^{2}\mathrm{d}t\right]+C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left\|{\Phi^{n}}\right\|^{10}\mathrm{d}t\right]+\frac{1}{4}\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t\right].$ (4.24)
For the term $J_{3}$, we use the Lipschitz assumptions on $g$ to obtain
$J_{3}$ $\leq C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left(1+\left\|{\Phi^{n}}\right\|^{2}\right)\mathrm{d}t\right].$ (4.25)
Finally, for the term $J_{4}$, we apply the Burkholder–Davis–Gundy inequality to have
$J_{4}$ $\leq C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left\|{\Phi^{n}}\right\|^{2}\sum_{k=1}^{\infty}\left\|{g_{k}(\Phi^{n})}\right\|^{2}\mathrm{d}t\right]^{\frac{1}{2}}$
$\leq\frac{1}{2}\mathbb{E}\left[\sup_{t\in[0,S\wedge\tau]}\left\|{\Phi^{n}}\right\|^{2}\right]+C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left(1+\left\|{\Phi^{n}}\right\|^{2}\right)\mathrm{d}t\right]$ (4.26)
Substituting (4.23), (4), (4.25), and (4) back into (4), then rearranging the terms, we obtain
$\mathbb{E}\left[\sup_{t\in[0,S\wedge\tau]}\left\|{\Phi^{n}}\right\|^{2}\right]+2\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t\right]$
$\leq\left\|{\Phi_{0}}\right\|^{2}+C\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left(1+\left\|{\Phi^{n}}\right\|^{2}+\left\|{\Phi^{n}}\right\|^{10}\right)\mathrm{d}t\right],$
where $C$ is independent of $\tau,S$, and $n$. Therefore, by applying Chebyshev’s inequality, for the set $A_{n}(\tau,S)$ defined in (4.6) we have for any $T>0$ and $M>1$,
$\mathbb{P}\left[A_{n}(\tau,S)\right]$ $\leq\mathbb{P}\left[C\int_{0}^{S\wedge\tau}\left(1+\left\|{\Phi^{n}}\right\|^{2}+\left\|{\Phi^{n}}\right\|^{10}\right)\mathrm{d}t>(M-1)^{2}\right]$
$\leq C(M-1)^{-2}\,\mathbb{E}\left[\int_{0}^{S\wedge\tau}\left(1+\left\|{\Phi^{n}}\right\|^{2}+\left\|{\Phi^{n}}\right\|^{10}\right)\mathrm{d}t\right]\leq CS,$
where $C=C(M)$, and in the last step we used the fact that $\tau\in\mathcal{T}_{n}^{M,T}$. Taking supremum over $\tau\in\mathcal{T}_{n}^{M,T}$, then over $n\in\mathbb{N}$, and taking the limit as $S\to 0$, we deduce (4.8). This completes the proof of the proposition. ∎
We now prove the existence of a local strong pathwise solution to (1.1).
Theorem 4.6.
Let $\Phi_{0}\in\mathcal{V}$ be a given initial data. Then there exists a local strong solution $(\Phi,\tau)$ to the problem (1.1) in the sense of Definition (4.2).
Proof.
We proceed with a compactness argument by applying Theorem A.1 with $X=\mathcal{V}$ and $Y=\mathrm{D}(\mathcal{A})$, noting the definition of $\mathcal{T}_{n}^{M,T}$ in (4.5). Since the hypotheses of this theorem have been shown to hold in Proposition 4.5, we infer the existence of a subsequence of the Galerkin solutions $\{\Phi^{n_{\ell}}\}_{\ell\in\mathbb{N}}$ to (4.4), a strictly positive stopping time $\tau\leq T$, and a process $\Phi(\cdot)=\Phi(\,\cdot\wedge\tau)$ which is continuous in $\mathcal{V}$, such that
$\lim_{\ell\to\infty}\left(\sup_{t\in[0,\tau]}\left\|{\Phi^{n_{\ell}}(t)-\Phi(t)}\right\|^{2}+\int_{0}^{\tau}\left|{\mathcal{A}\big(\Phi^{n_{\ell}}(t)-\Phi(t)\big)}\right|^{2}\mathrm{d}t\right)=0\;\text{ a.s.}$ (4.27)
Furthermore, (A.4) implies for any $p\in[1,\infty)$,
$\Phi(\,\cdot\wedge\tau)\in L^{p}\big(\Omega;C([0,T];\mathcal{V})\big)\;\text{ and }\;\Phi\mathds{1}_{t\leq\tau}\in L^{p}\big(\Omega;L^{2}(0,T;\mathrm{D}(\mathcal{A}))\big),$ (4.28)
while (A.5) confers the existence of a collection of measurable sets $\{\Omega_{\ell}\}_{\ell\in\mathbb{N}}$ with
$\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\left(\sup_{t\in[0,\tau]}\left\|{\Phi^{n_{\ell}}(t)}\right\|^{2}+\int_{0}^{\tau}\left|{\mathcal{A}\Phi^{n_{\ell}}(t)}\right|^{2}\mathrm{d}t\right)^{\frac{p}{2}}\right]<\infty.$ (4.29)
Thus, given (4.27) and (4.29), by Lemma A.2 we have
$\mathds{1}_{t\leq\tau}\mathds{1}_{\Omega_{\ell}}\Phi^{n_{\ell}}\rightharpoonup\mathds{1}_{t\leq\tau}\Phi\quad\text{in }L^{p}\big(\Omega;L^{2}(0,T;\mathrm{D}(\mathcal{A}))\big),$ (4.30)
$\mathds{1}_{\Omega_{\ell}}\Phi^{n_{\ell}}(\,\cdot\wedge\tau)\overset{\ast}{\rightharpoonup}\Phi\quad\text{in }L^{p}\big(\Omega;L^{\infty}(0,T;\mathcal{V})\big).$ (4.31)
Next, we consider the convergence of the first nonlinear term $P_{n_{\ell}}\mathcal{B}(\Phi^{n_{\ell}})$. For any $v\in\mathcal{V}$, we have
$\left|{\left\langle P_{n_{\ell}}\mathcal{B}(\Phi^{n_{\ell}})-\mathcal{B}(\Phi),v\right\rangle}\right|$ $\leq\left|{\left\langle\mathcal{B}(\Phi^{n_{\ell}}-\Phi,\Phi^{n_{\ell}}),P_{n_{\ell}}v\right\rangle}\right|+\left|{\left\langle\mathcal{B}(\Phi,\Phi^{n_{\ell}}-\Phi),P_{n_{\ell}}v\right\rangle}\right|+\left|{\left\langle\mathcal{B}(\Phi),Q_{n_{\ell}}v\right\rangle}\right|$
$\leq C\left\|{\Phi^{n_{\ell}}-\Phi}\right\|\left\|{\Phi^{n_{\ell}}}\right\|\left\|{v}\right\|+C\left\|{\Phi}\right\|\left\|{\Phi^{n_{\ell}}-\Phi}\right\|\left\|{v}\right\|+C\left\|{\Phi}\right\|^{2}\left|{Q_{n_{\ell}}v}\right|^{\frac{1}{2}}\left\|{Q_{n_{\ell}}v}\right\|^{\frac{1}{2}}$
$\leq C\left\|{\Phi^{n_{\ell}}-\Phi}\right\|\left\|{v}\right\|\left(\left\|{\Phi^{n_{\ell}}}\right\|+\left\|{\Phi}\right\|\right)+C\lambda_{n_{\ell}}^{-\frac{1}{4}}\left\|{\Phi}\right\|^{2}\left\|{v}\right\|,$
where in the penultimate step we used (2.2), while in the last step we applied (A.8). This estimate, together with (4.27), implies that given any $v\in\mathcal{V}$, for a.e. $(\omega,t)\in\Omega\times[0,T]$,
[…]
Furthermore, we have
$\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\int_{0}^{\tau}\left|{P_{n_{\ell}}\mathcal{B}(\Phi^{n_{\ell}})}\right|^{2}\mathrm{d}t\right]$ $\leq C\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\int_{0}^{\tau}\left\|{\Phi^{n_{\ell}}}\right\|^{3}\left|{\mathcal{A}\Phi^{n_{\ell}}}\right|\mathrm{d}t\right]$
$\leq C\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\left(\sup_{s\in[0,\tau]}\left\|{\Phi^{n_{\ell}}(s)}\right\|^{2}\right)\left(\int_{0}^{\tau}\left|{\mathcal{A}\Phi^{n_{\ell}}}\right|^{2}\mathrm{d}t\right)\right]$
$\leq C\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\left(\sup_{s\in[0,\tau]}\left\|{\Phi^{n_{\ell}}(s)}\right\|^{4}+\left(\int_{0}^{\tau}\left|{\mathcal{A}\Phi^{n_{\ell}}}\right|^{2}\mathrm{d}t\right)^{2}\right)\right]<\infty,$ (4.33)
where we applied Young’s inequality and (4.29) in the last line. Altogether, (4.32), (4), and Lemma A.2 imply that
$\mathds{1}_{t\leq\tau}\mathds{1}_{\Omega_{\ell}}P_{n_{\ell}}\mathcal{B}(\Phi^{n_{\ell}})\rightharpoonup\mathds{1}_{t\leq\tau}\mathcal{B}(\Phi)\quad\text{in }L^{2}\big(\Omega;L^{2}(0,T;\mathcal{H})\big).$ (4.34)
Now, we study the convergence of the second nonlinear term $P_{n_{\ell}}\mathcal{R}(\Phi^{n_{\ell}})$. For any $v\in\mathcal{V}$, by the triangle and the Hölder inequalities we have
$\left|{\left\langle P_{n_{\ell}}\mathcal{R}(\Phi^{n_{\ell}})-\mathcal{R}(\Phi),v\right\rangle}\right|$ $\leq\left|{\left\langle\mathcal{R}(\Phi^{n_{\ell}})-\mathcal{R}(\Phi),P_{n_{\ell}}v\right\rangle}\right|+\left|{\left\langle\mathcal{R}(\Phi),Q_{n_{\ell}}v\right\rangle}\right|$
$\leq C\left(1+\left\|{\Phi^{n_{\ell}}}\right\|^{2}+\left\|{\Phi}\right\|^{2}\right)\left\|{\Phi^{n_{\ell}}-\Phi}\right\|\left|{v}\right|+C\left(1+\left\|{\Phi}\right\|^{2}\right)\left\|{\Phi}\right\|\left|{Q_{n_{\ell}}v}\right|$
$\leq C\left(1+\left\|{\Phi^{n_{\ell}}}\right\|^{2}+\left\|{\Phi}\right\|^{2}\right)\left\|{\Phi^{n_{\ell}}-\Phi}\right\|\left\|{v}\right\|+C\lambda_{n_{\ell}}^{-\frac{1}{2}}\left(1+\left\|{\Phi}\right\|^{2}\right)\left\|{\Phi}\right\|\left\|{v}\right\|.$
Here, in the penultimate step we utilised (2.9) and (2.7), while in the last step we applied (A.8). Similarly to (4.32), the above estimate, together with (4.27), implies that for a.e. $(\omega,t)\in\Omega\times[0,T]$,
[…]
Moreover, noting (4.29), we have the estimate
$\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\int_{0}^{\tau}\left|{P_{n_{\ell}}\mathcal{R}(\Phi^{n_{\ell}})}\right|^{2}\mathrm{d}t\right]$ $\leq C\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\int_{0}^{\tau}\left(1+\left\|{\Phi^{n_{\ell}}}\right\|^{4}\right)\left\|{\Phi^{n_{\ell}}}\right\|^{2}\mathrm{d}t\right]$
$\leq C\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\left(\sup_{s\in[0,\tau]}\left\|{\Phi^{n_{\ell}}(s)}\right\|^{2}+\left\|{\Phi^{n_{\ell}}(s)}\right\|^{6}\right)\right]<\infty.$ (4.36)
Thus, (4.35), (4), and Lemma A.2 imply that
$\mathds{1}_{t\leq\tau}\mathds{1}_{\Omega_{\ell}}P_{n_{\ell}}\mathcal{R}(\Phi^{n_{\ell}})\rightharpoonup\mathds{1}_{t\leq\tau}\mathcal{R}(\Phi)\quad\text{in }L^{2}\big(\Omega;L^{2}(0,T;\mathcal{H})\big).$ (4.37)
Finally, we consider the stochastic term. Noting the assumptions on $g$, we obtain
$\left(\sum_{k=1}^{\infty}\left|{P_{n_{\ell}}g_{k}(\Phi^{n_{\ell}})-g_{k}(\Phi)}\right|^{2}\right)^{\frac{1}{2}}$ $\leq\left(\sum_{k=1}^{\infty}\left|{P_{n_{\ell}}g_{k}(\Phi^{n_{\ell}})-P_{n_{\ell}}g_{k}(\Phi)}\right|^{2}\right)^{\frac{1}{2}}+\left(\sum_{k=1}^{\infty}\left|{Q_{n_{\ell}}g_{k}(\Phi)}\right|^{2}\right)^{\frac{1}{2}}$
$\leq C\left(\left\|{\Phi^{n_{\ell}}-\Phi}\right\|+\lambda_{n_{\ell}}^{-1}(1+\left\|{\Phi}\right\|)\right).$
This implies for a.e. $(\omega,t)\in\Omega\times[0,T]$, we have $\mathds{1}_{t\leq\tau}P_{n_{\ell}}g(\Phi^{n_{\ell}})\to\mathds{1}_{t\leq\tau}g(\Phi)$ in $\ell^{2}(\mathcal{H})$. Furthermore, again by the assumptions on $g$ and (4.29),
$\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\int_{0}^{\tau}\sum_{k=1}^{\infty}\left|{P_{n_{\ell}}g_{k}(\Phi^{n_{\ell}})}\right|^{2}\mathrm{d}t\right]\leq C+C\sup_{\ell\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{\ell}}\int_{0}^{\tau}\left\|{\Phi^{n_{\ell}}}\right\|^{2}\mathrm{d}t\right]<\infty,$
from which we infer that
$\mathds{1}_{t\leq\tau}\mathds{1}_{\Omega_{\ell}}P_{n_{\ell}}g(\Phi^{n_{\ell}})\rightharpoonup\mathds{1}_{t\leq\tau}g(\Phi)\quad\text{in }L^{2}\big(\Omega;L^{2}(0,T;\ell^{2}(\mathcal{H}))\big).$ (4.38)
Thus, from (4.30), (4.34), (4.37), and (4.38), we deduce for any fixed $v\in\mathcal{H}$ the following weak convergences in $L^{2}\big(\Omega;L^{2}(0,T)\big)$:
$\mathds{1}_{\Omega_{\ell}}\int_{0}^{t\wedge\tau}\left\langle\mathcal{A}\Phi^{n_{\ell}},v\right\rangle\mathrm{d}s$ $\rightharpoonup\int_{0}^{t\wedge\tau}\left\langle\mathcal{A}\Phi,v\right\rangle\mathrm{d}s,$ (4.39)
$\mathds{1}_{\Omega_{\ell}}\int_{0}^{t\wedge\tau}\left\langle P_{n_{\ell}}\mathcal{B}(\Phi^{n_{\ell}}),v\right\rangle\mathrm{d}s$ $\rightharpoonup\int_{0}^{t\wedge\tau}\left\langle\mathcal{B}(\Phi),v\right\rangle\mathrm{d}s,$
$\mathds{1}_{\Omega_{\ell}}\sum_{k=1}^{\infty}\int_{0}^{t\wedge\tau}\left\langle P_{n_{\ell}}g_{k}(\Phi^{n_{\ell}}),v\right\rangle\mathrm{d}\beta_{k}$ $\rightharpoonup\sum_{k=1}^{\infty}\int_{0}^{t\wedge\tau}\left\langle g_{k}(\Phi),v\right\rangle\mathrm{d}\beta_{k}.$
Finally, we apply the following standard argument. Let $K\subset\Omega\times[0,T]$ be a measurable set. Then by (4.31) and (4.39), we have
$\mathbb{E}\left[\int_{0}^{T}\mathds{1}_{K}\left\langle\Phi(t),v\right\rangle\mathrm{d}t\right]$ $=\lim_{\ell\to\infty}\mathbb{E}\left[\int_{0}^{T}\left\langle\mathds{1}_{\Omega_{\ell}}\Phi^{n_{\ell}}(t\wedge\tau),\mathds{1}_{K}v\right\rangle\mathrm{d}t\right]$
$=\lim_{\ell\to\infty}\mathbb{E}\left[\int_{0}^{T}\mathds{1}_{K}\mathds{1}_{\Omega_{\ell}}\left\langle P_{n_{\ell}}\Phi_{0},v\right\rangle\mathrm{d}t\right]$
$\quad-\lim_{\ell\to\infty}\mathbb{E}\left[\int_{0}^{T}\mathds{1}_{K}\mathds{1}_{\Omega_{\ell}}\left(\int_{0}^{t\wedge\tau}\left\langle\mathcal{A}\Phi^{n_{\ell}}+P_{n_{\ell}}\mathcal{B}(\Phi^{n_{\ell}})+P_{n_{\ell}}\mathcal{R}(\Phi^{n_{\ell}}),v\right\rangle\mathrm{d}s\right)\mathrm{d}t\right]$
$\quad+\lim_{\ell\to\infty}\mathbb{E}\left[\int_{0}^{T}\mathds{1}_{K}\mathds{1}_{\Omega_{\ell}}\left(\sum_{k=1}^{\infty}\int_{0}^{t\wedge\tau}\left\langle P_{n_{\ell}}g_{k}(\Phi^{n_{\ell}}),v\right\rangle\mathrm{d}\beta_{k}\right)\mathrm{d}t\right]$
$=\mathbb{E}\left[\int_{0}^{T}\mathds{1}_{K}\left(\left\langle\Phi_{0},v\right\rangle-\int_{0}^{t\wedge\tau}\left\langle\mathcal{A}\Phi+\mathcal{B}(\Phi)+\mathcal{R}(\Phi),v\right\rangle\mathrm{d}s\right)\mathrm{d}t\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\mathds{1}_{K}\left(\sum_{k=1}^{\infty}\int_{0}^{t\wedge\tau}\left\langle g_{k}(\Phi),v\right\rangle\mathrm{d}\beta_{k}\right)\mathrm{d}t\right].$
Since $v$ and $K$ are arbitrary, we infer that $\Phi$ satisfies (4.2) as an equation in $\mathcal{H}$. The required regularity for a strong solution is conferred by (4.28). This completes the proof of the theorem. ∎
The following theorem establishes the pathwise uniqueness of local strong solution obtained previously.
Theorem 4.7.
Let $\tau>0$ be a stopping time such that $(\Phi^{(1)},\tau)$ and $(\Phi^{(2)},\tau)$ are local strong solutions to (1.1) with the same initial data $\Phi_{0}$. Then
$\mathbb{P}\left[\Phi^{(1)}(t\wedge\tau)=\Phi^{(2)}(t\wedge\tau),\;\forall t\in[0,\infty)\right]=1.$ (4.40)
Proof.
Let $\Psi:=\Phi^{(1)}-\Phi^{(2)}$. Then $\Psi$ satisfies
$\mathrm{d}\Psi+\left(\mathcal{A}\Psi+\mathcal{B}(\Phi^{(1)})-\mathcal{B}(\Phi^{(2)})+\mathcal{R}(\Phi^{(1)})-\mathcal{R}(\Phi^{(2)})\right)\mathrm{d}t=\sum_{k=1}^{\infty}\left(g_{k}(\Phi^{(1)})-g_{k}(\Phi^{(2)})\right)\mathrm{d}\beta_{k}.$
Applying Itô’s lemma to $\left|{\Psi}\right|^{2}$ yields
$\mathrm{d}\left|{\Psi}\right|^{2}$ $=-2\left\|{\Psi}\right\|^{2}\mathrm{d}t-2\left\langle\mathcal{B}(\Phi^{(1)})-\mathcal{B}(\Phi^{(2)}),\Psi\right\rangle\mathrm{d}t-2\left\langle\mathcal{R}(\Phi^{(1)})-\mathcal{R}(\Phi^{(2)}),\Psi\right\rangle\mathrm{d}t$
$\quad+\sum_{k=1}^{\infty}\left|{g_{k}(\Phi^{(1)})-g_{k}(\Phi^{(2)})}\right|^{2}\mathrm{d}t+2\sum_{k=1}^{\infty}\left\langle g_{k}(\Phi^{(1)})-g_{k}(\Phi^{(2)}),\Psi\right\rangle\mathrm{d}\beta_{k}.$
For any stopping time $\sigma\leq\tau$, we integrate the above expression over $0,\sigma]$ and apply the expected value to obtain
$\mathbb{E}\left[\left|{\Psi(\sigma)}\right|^{2}\right]+2\mathbb{E}\left[\int_{0}^{\sigma}\left\|{\Psi(t)}\right\|^{2}\mathrm{d}t\right]$
$=-2\mathbb{E}\left[\int_{0}^{\sigma}\left\langle\mathcal{B}(\Phi^{(1)})-\mathcal{B}(\Phi^{(2)}),\Psi\right\rangle\mathrm{d}t\right]-2\mathbb{E}\left[\int_{0}^{\sigma}\left\langle\mathcal{R}(\Phi^{(1)})-\mathcal{R}(\Phi^{(2)}),\Psi\right\rangle\mathrm{d}t\right]$
$\quad+\mathbb{E}\left[\int_{0}^{\sigma}\left|{g_{k}(\Phi^{(1)})-g_{k}(\Phi^{(2)})}\right|^{2}\mathrm{d}t\right],$ (4.41)
where we also used the martingale property of stochastic integral. Note that we have the following estimates. Firstly, by (2.1) and (2.2),
$\left|{\left\langle\mathcal{B}(\Phi^{(1)})-\mathcal{B}(\Phi^{(2)}),\Psi\right\rangle}\right|=\left|{\left\langle\mathcal{B}(\Psi,\Phi^{(1)}),\Psi\right\rangle}\right|$ $\leq C\left\|{\Psi}\right\|^{\frac{3}{2}}\left\|{\Phi^{(1)}}\right\|\left|{\Psi}\right|^{\frac{1}{2}}$
$\leq\frac{1}{4}\left\|{\Psi}\right\|^{2}+C\left\|{\Phi^{(1)}}\right\|^{4}\left|{\Psi}\right|^{2},$ (4.42)
where in the last step we used Young’s inequality. Furthermore, by (2.9), we have
$\left|{\left\langle\mathcal{R}(\Phi^{(1)})-\mathcal{R}(\Phi^{(2)}),\Psi\right\rangle}\right|$ $=C\left(1+\left\|{\Phi^{(1)}}\right\|^{2}+\left\|{\Phi^{(2)}}\right\|^{2}\right)\left\|{\Psi}\right\|\left|{\Psi}\right|$
$\leq\frac{1}{4}\left\|{\Psi}\right\|^{2}+C\left(1+\left\|{\Phi^{(1)}}\right\|^{4}+\left\|{\Phi^{(2)}}\right\|^{4}\right)\left|{\Psi}\right|^{2}.$ (4.43)
Now, for any $R>0$, define the stopping time
$\sigma_{R}:=\inf\left\{t>0:\left\|{\Phi^{(1)}(t)}\right\|^{2}+\left\|{\Phi^{(2)}(t)}\right\|^{2}>R\right\}\wedge\tau.$
Noting this stopping time and Lipschitz assumptions on $g$, we employ estimates (4) and (4) in (4) and rearrange the terms to obtain
$\mathbb{E}\left[\left|{\Psi(\sigma_{R}\wedge t)}\right|^{2}\right]+\mathbb{E}\left[\int_{0}^{\sigma_{R}\wedge t}\left\|{\Psi(s)}\right\|^{2}\mathrm{d}s\right]$ $\leq C\mathbb{E}\left[\int_{0}^{\sigma_{R}\wedge t}\left(1+\left\|{\Phi^{(1)}}\right\|^{4}+\left\|{\Phi^{(2)}}\right\|^{4}\right)\left|{\Psi(s)}\right|^{2}\mathrm{d}s\right]$
$\leq C(1+R^{2})\,\mathbb{E}\left[\int_{0}^{\sigma_{R}\wedge t}\left|{\Psi(s)}\right|^{2}\mathrm{d}s\right].$
By the Gronwall lemma, we infer that
$\mathbb{E}\left[\left|{\Psi(\sigma_{R}\wedge t)}\right|^{2}\right]=0,$
which implies $\Phi^{(1)}(\sigma_{R}\wedge t)=\Phi^{(2)}(\sigma_{R}\wedge t)$ a.s. Thus,
$\mathbb{P}\left[\Phi^{(1)}(t\wedge\tau)\neq\Phi^{(2)}(t\wedge\tau)\right]$ $\leq\mathbb{P}\left[\{\sigma_{R}<\tau\}\cap\{\Phi^{(1)}(t\wedge\tau)\neq\Phi^{(2)}(t\wedge\tau)\}\right]$
$\leq\mathbb{P}\left[\sigma_{R}<\tau\right]$
$\leq\mathbb{P}\left[\sup_{s\in[0,\tau]}\left(\left\|{\Phi^{(1)}(s)}\right\|^{2}+\left\|{\Phi^{(2)}(s)}\right\|^{2}\right)>R\right]$
$\leq\frac{1}{R}\mathbb{E}\left[\sup_{s\in[0,\tau]}\left(\left\|{\Phi^{(1)}(s)}\right\|^{2}+\left\|{\Phi^{(2)}(s)}\right\|^{2}\right)\right],$
which tends to 0 as $R\to\infty$. Therefore for any $t\geq 0$, we have $\Phi^{(1)}(t\wedge\tau)=\Phi^{(2)}(t\wedge\tau)$ on a set of full measure which may depend on $t$. By taking the intersection of such sets corresponding to positive rational times, we obtain
$\mathbb{P}\left[\Phi^{(1)}(t\wedge\tau)=\Phi^{(2)}(t\wedge\tau),\;\forall t\in[0,\infty)\cap\mathbb{Q}\right]=1.$
By the continuity in time property (4.3) of the solution, we deduce (4.40), completing the proof of the theorem. ∎
5. Existence of a maximal strong solution
In this section, we establish the existence of a maximal strong solution $(\Phi,\xi)$ to (1.1). Furthermore, we show that this maximal strong solution is global if $d=2$. We begin with the following estimate on a local weak solution of (1.1) which will be used on several occasions.
Lemma 5.1.
Suppose that $(\Phi,\tau)$ is a local weak solution of (1.1) in the sense of Definition 4.1 corresponding to an initial data $\Phi_{0}\in\mathcal{H}$. Then for any $T>0$ and $p\geq 4$,
$\mathbb{E}\left[\sup_{t\in[0,\tau\wedge T]}\left|{\Phi(t)}\right|^{p}\right]+\mathbb{E}\left[\int_{0}^{\tau\wedge T}\left\|{\Phi(s)}\right\|^{2}\left|{\Phi(s)}\right|^{p-2}\mathrm{d}s\right]\leq C_{p}\left(1+\left|{\Phi_{0}}\right|^{p}\right),$
where $C_{p}$ is a constant which depends on $p$ and $T$, but is independent of $\tau$.
Proof.
For any stopping times $\sigma_{a},\sigma_{b}$ such that $0\leq\sigma_{a}\leq t\leq\sigma_{b}\leq\tau\wedge T$, by applying Itô’s lemma to $\left|{\Phi}\right|^{p}$, where $p\geq 4$, we have
$\left|{\Phi(t)}\right|^{p}+p\int_{\sigma_{a}}^{t}\left\|{\Phi(s)}\right\|^{2}\left|{\Phi(s)}\right|^{p-2}\mathrm{d}s+p\int_{\sigma_{a}}^{t}\left|{\Phi(s)}\right|^{p-2}\left\langle\mathcal{R}(\Phi(s)),\Phi(s)\right\rangle\mathrm{d}s$
$=\left|{\Phi(\sigma_{a})}\right|^{p}+\frac{p}{2}\int_{\sigma_{a}}^{t}\sum_{k=1}^{\infty}\left|{\Phi(s)}\right|^{p-2}\left|{g_{k}(\Phi(s))}\right|^{2}\mathrm{d}s+\frac{p(p-2)}{2}\int_{\sigma_{a}}^{t}\sum_{k=1}^{\infty}\left|{\left\langle g_{k}(\Phi(s)),\Phi(s)\right\rangle}\right|^{2}\left|{\Phi(s)}\right|^{p-4}\mathrm{d}s$
$\quad+p\int_{\sigma_{a}}^{t}\sum_{k=1}^{\infty}\left|{\Phi(s)}\right|^{p-2}\left\langle g_{k}(\Phi(s)),\Phi(s)\right\rangle\mathrm{d}\beta_{k}.$
Noting assumption (2.6) on $\mathcal{R}$, we then have
$\left|{\Phi(t)}\right|^{p}+p\int_{\sigma_{a}}^{t}\left\|{\Phi(s)}\right\|^{2}\left|{\Phi(s)}\right|^{p-2}\mathrm{d}s$
$\leq\left|{\Phi(\sigma_{a})}\right|^{p}+p\int_{\sigma_{a}}^{t}\left|{\Phi(s)}\right|^{p-2}\left|{\left\langle\mathcal{F}(\Phi(s)),\Phi(s)\right\rangle}\right|\mathrm{d}s+\frac{p(p-1)}{2}\int_{\sigma_{a}}^{t}\sum_{k=1}^{\infty}\left|{\Phi(s)}\right|^{p-2}\left|{g_{k}(\Phi(s))}\right|^{2}\mathrm{d}s$
$\quad+p\int_{\sigma_{a}}^{t}\sum_{k=1}^{\infty}\left|{\Phi(s)}\right|^{p-2}\left\langle g_{k}(\Phi(s)),\Phi(s)\right\rangle\mathrm{d}\beta_{k}.$
Using the triangle inequality, taking supremum over $t\in[\sigma_{a},\sigma_{b}]$, and applying the expected value, we obtain
$\mathbb{E}\left[\sup_{t\in[\sigma_{a},\sigma_{b}]}\left|{\Phi(t)}\right|^{p}\right]+p\mathbb{E}\left[\int_{\sigma_{a}}^{\sigma_{b}}\left\|{\Phi(s)}\right\|^{2}\left|{\Phi(s)}\right|^{p-2}\mathrm{d}s\right]$
$\leq C_{p}\mathbb{E}\left[\left|{\Phi(\sigma_{a})}\right|^{p}\right]+C_{p}\mathbb{E}\left[\int_{\sigma_{a}}^{\sigma_{b}}\left|{\Phi(s)}\right|^{p-2}\left|{\left\langle\mathcal{F}(\Phi(s)),\Phi(s)\right\rangle}\right|\mathrm{d}s\right]$
$\quad+C_{p}\mathbb{E}\left[\int_{\sigma_{a}}^{\sigma_{b}}\sum_{k=1}^{\infty}\left|{\Phi(s)}\right|^{p-2}\left|{g_{k}(\Phi(s))}\right|^{2}\mathrm{d}s\right]$
$\quad+C_{p}\mathbb{E}\left[\sup_{t\in[\sigma_{a},\sigma_{b}]}\left|{\int_{\sigma_{a}}^{t}\sum_{k=1}^{\infty}\left|{\Phi(s)}\right|^{p-2}\left\langle g_{k}(\Phi(s)),\Phi(s)\right\rangle\mathrm{d}\beta_{k}}\right|\right]$
$=:C_{p}\mathbb{E}\left[\left|{\Phi(\sigma_{a})}\right|^{p}\right]+I_{1}+I_{2}+I_{3}.$ (5.1)
We will estimate each term on the last line. Firstly, note that by (2.6), we have
$\left|{\Phi(s)}\right|^{p-2}\left|{\left\langle\mathcal{F}(\Phi(s)),\Phi(s)\right\rangle}\right|$ $\leq C\left|{\Phi(s)}\right|^{p-2}\left(1+\left\|{\Phi(s)}\right\|\right)\left|{\Phi(s)}\right|$
$\leq C\left|{\Phi(s)}\right|^{p-1}+C\left|{\Phi(s)}\right|^{\frac{p-2}{2}}\left\|{\Phi(s)}\right\|\left|{\Phi(s)}\right|^{\frac{p}{2}}$
$\leq C+C\left|{\Phi(s)}\right|^{p}+\frac{p}{4}\left\|{\Phi(s)}\right\|^{2}\left|{\Phi(s)}\right|^{p-2},$ (5.2)
where in the last step we applied Young’s inequality. Next, by the assumptions on $g$ and Young’s inequality, we obtain
$\int_{\sigma_{a}}^{\sigma_{b}}\sum_{k=1}^{\infty}\left|{\Phi(s)}\right|^{p-2}\left|{g_{k}(\Phi(s))}\right|^{2}\mathrm{d}s$ $\leq C\int_{\sigma_{a}}^{\sigma_{b}}\left|{\Phi(s)}\right|^{p-2}\left(1+\left|{\Phi(s)}\right|^{2}\right)\mathrm{d}s$
$\leq C\left(1+\int_{\sigma_{a}}^{\sigma_{b}}\left|{\Phi(s)}\right|^{p}\mathrm{d}s\right).$ (5.3)
Similarly, by the Burkholder–Davis–Gundy inequality, we also have
$\mathbb{E}\left[\sup_{t\in[\sigma_{a},\sigma_{b}]}\left|{\sum_{k=1}^{\infty}\int_{\sigma_{a}}^{t}\left|{\Phi(s)}\right|^{p-2}\left\langle g_{k}(\Phi(s)),\Phi(s)\right\rangle\mathrm{d}\beta_{k}}\right|\right]$
$\leq C_{p}\mathbb{E}\left[\int_{\sigma_{a}}^{\sigma_{b}}\sum_{k=1}^{\infty}\left|{\left\langle g_{k}(\Phi(s)),\Phi(s)\right\rangle}\right|^{2}\left|{\Phi(s)}\right|^{2(p-2)}\mathrm{d}s\right]^{\frac{1}{2}}$
$\leq C_{p}\mathbb{E}\left[\int_{\sigma_{a}}^{\sigma_{b}}\left(1+\left|{\Phi(s)}\right|^{2}\right)\left|{\Phi(s)}\right|^{2(p-1)}\mathrm{d}s\right]^{\frac{1}{2}}$
$\leq\frac{1}{4}\mathbb{E}\left[\sup_{t\in[\sigma_{a},\sigma_{b}]}\left|{\Phi(t)}\right|^{p}\right]+C\mathbb{E}\left[1+\int_{\sigma_{a}}^{\sigma_{b}}\left|{\Phi(s)}\right|^{p}\mathrm{d}s\right],$ (5.4)
where in the last step we again used Young’s inequality. Altogether, applying (5), (5), and (5) in (5), and rearranging the terms, we infer
$\mathbb{E}\left[\sup_{t\in[\sigma_{a},\sigma_{b}]}\left|{\Phi(t)}\right|^{p}\right]+p\mathbb{E}\left[\int_{\sigma_{a}}^{\sigma_{b}}\left\|{\Phi(s)}\right\|^{2}\left|{\Phi(s)}\right|^{p-2}\mathrm{d}s\right]\leq\mathbb{E}\left[1+\left|{\Phi(\sigma_{a})}\right|^{p}\right]+C_{p}\mathbb{E}\left[\int_{\sigma_{a}}^{\sigma_{b}}\left|{\Phi(s)}\right|^{p}\mathrm{d}s\right].$
The required inequality then follows from Lemma A.4. ∎
5.1. Maximal strong solution for $d\leq 3$
The existence of a maximal strong pathwise solution in the sense of Definition 4.3 will be shown in this section. The argument follows closely the idea in [7, Section 5.3.4]
Theorem 5.2.
There exists a unique maximal strong pathwise solution $(\Phi,\xi)$ to (1.1) and a sequence of stopping times $\{\tau_{n}\}_{n\in\mathbb{N}}$ announcing $\xi$.
Proof.
With the aim of extending the solution to a maximal time of existence, we define the set
[…]
The set $\mathcal{L}$ is non-empty by Theorem 4.6. More precisely, every stopping time $\tau_{n}$ defined in Remark 4.4 belongs to $\mathcal{L}$ and $\xi=\sup_{n}\tau_{n}$ is a stopping time. For such $\tau_{n}$ and each $n\in\mathbb{N}$, let $\Phi_{n}$ denote the process for which $(\Phi_{n},\tau_{n})$ is a local pathwise strong solution. For $\tau_{k}<\tau_{l}$ define
$\Omega_{k,\ell}:=\{\omega\in\Omega:\Phi_{k}(t\wedge\tau_{k})=\Phi_{\ell}(t\wedge\tau_{k}),\;\forall t\in[0,\infty)\}$
By uniqueness (Theorem 4.7), we have $\mathbb{P}\big[\widetilde{\Omega}\big]=1$, where $\widetilde{\Omega}:=\bigcap_{k,\ell\in\mathbb{N}}\Omega_{k,\ell}$. For a fixed $\omega\in\widetilde{\Omega}$ and every $t>0$, the sequence $\{\Phi_{k}(t\wedge\tau_{k})\}_{k\in\mathbb{N}}$ is Cauchy in $\mathcal{V}$. Let
$\widetilde{\Phi}(t):=\lim_{k\to\infty}\Phi_{k}(t\wedge\tau_{k})\,\text{ a.s.}$
By Lemma 5.1 and the monotone convergence theorem, we have for any $T>0$,
$\mathbb{E}\left[\sup_{t\in[0,\xi\wedge T]}\left|{\widetilde{\Phi}(t)}\right|^{2}+\int_{0}^{\xi\wedge T}\left\|{\widetilde{\Phi}(s)}\right\|^{2}\mathrm{d}s\right]<\infty.$ (5.5)
Therefore, noting (5.5), we can define $\Phi(t)$ for any $t>0$ by
$\left\langle\Phi(t),v\right\rangle$ $=\left\langle\Phi(0),v\right\rangle-\int_{0}^{t\wedge\xi}\left\langle\mathcal{A}\widetilde{\Phi}+\mathcal{B}(\widetilde{\Phi})-\mathcal{R}(\widetilde{\Phi}),v\right\rangle\mathrm{d}s+\int_{0}^{t\wedge\xi}\sum_{k=1}^{\infty}\left\langle g_{k}(\widetilde{\Phi}),v\right\rangle\mathrm{d}\beta_{k}$ (5.6)
for all $v\in\mathcal{V}$. Note that $\Phi(t,\omega)=\widetilde{\Phi}(t,\omega)$ for all $t<\xi$ and $\Phi(\cdot)$ is weakly continuous a.s. in $\mathcal{H}$. As such, (5.6) implies that $(\Phi,\xi)$ is a local weak pathwise solution in the sense of Definition 4.1.
[…]
$\Phi_{\mathrm{e}}(t):=\begin{cases}\Phi(t),&\text{if $t\leq\tau_{n}$,}\\ \hat{\Phi}(t),&\text{if $\tau_{n}<t\leq\tau_{n}+\hat{\tau}$.}\end{cases}$
yields a local strong pathwise solution $(\Phi_{\mathrm{e}},\sigma)$, defined up to the stopping time $\sigma:=\tau_{n}+\hat{\tau}$, relative to the original stochastic basis $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$.
Finally, it remains to show that $\mathbb{P}\left[\tau_{n}\wedge T=\xi\right]=0$. Suppose otherwise, that for some sufficiently large $n$ and $T$, we have $\mathbb{P}\left[\tau_{n}\wedge T=\xi\right]>0$. The argument in the previous paragraph then confers the existence of a stopping time $\sigma\in\mathcal{L}$ such that $\sigma>\tau_{n}\wedge T$ a.s., contradicting the maximality of $\xi$ by the definition of $\xi=\sup\mathcal{L}$. We also note that $\{\tau_{n}\}_{n\in\mathbb{N}}$ announces $\xi$. This completes the proof of the theorem. ∎
5.2. Global strong solution for $d=2$
The existence of a global strong pathwise solution to (1.1) for $d=2$ will be established in this section.
Theorem 5.3.
Let $d=2$. The maximal strong pathwise solution $(\Phi,\xi)$ of (1.1) corresponding to an initial data $\Phi_{0}\in\mathcal{V}$ is global in the sense of Definition 4.3. In particular, there exists an increasing sequence of strictly positive stopping times $\{\tau_{n}\}_{n\in\mathbb{N}}$ with $\tau_{n}\uparrow\xi$, with $\mathbb{P}\left[\xi=\infty\right]=1$.
Proof.
Let $\{\tau_{n}\}_{n\in\mathbb{N}}$ be a strictly increasing sequence of stopping times announcing $\xi$. Note that
$\{\xi<\infty\}=\bigcup_{T=1}^{\infty}\{\xi\leq T\}=\bigcup_{T=1}^{\infty}\bigcap_{n=1}^{\infty}\{\tau_{n}\leq T\}.$ (5.7)
Since $\tau_{n}$ is increasing, we have
$\mathbb{P}\left[\bigcap_{n=1}^{\infty}\{\tau_{n}\leq T\}\right]=\lim_{n\to\infty}\mathbb{P}\left[\tau_{n}\leq T\right].$ (5.8)
Therefore, to show that $\mathbb{P}\left[\xi<\infty\right]=0$, by (5.7) and (5.8), it suffices to show that for any $T\geq 1$,
$\lim_{n\to\infty}\mathbb{P}\left[\tau_{n}\leq T\right]=0.$ (5.9)
To this end, for any $M>0$, define the stopping time
$\sigma_{M}:=\inf_{t\geq 0}\left\{\int_{0}^{t\wedge\xi}\left|{\Phi(s)}\right|^{6}\left\|{\Phi(s)}\right\|^{2}\mathrm{d}s>M\right\}\wedge 2T.$ (5.10)
Note that
$\mathbb{P}\left[\tau_{n}\leq T\right]\leq\mathbb{P}\left[\sigma_{M}\leq T\right]+\mathbb{P}\left[\left\{\tau_{n}\leq T\right\}\cap\left\{\sigma_{M}>T\right\}\right]=:I_{1}+I_{2}.$ (5.11)
The term $I_{1}$ can be bounded by the Chebyshev inequality as follows:
$I_{1}$ $\leq\mathbb{P}\left[\int_{0}^{T\wedge\xi}\left|{\Phi(s)}\right|^{6}\left\|{\Phi(s)}\right\|^{2}\mathrm{d}s\geq M\right]\leq\frac{1}{M}\mathbb{E}\left[\int_{0}^{T\wedge\xi}\left|{\Phi(s)}\right|^{6}\left\|{\Phi(s)}\right\|^{2}\mathrm{d}s\right].$ (5.12)
Note that the last term tends to zero as $M\to\infty$ by Lemma 5.1. Next, by definition, for the term $I_{2}$ in (5.11), we have
$I_{2}$ $\leq\mathbb{P}\left[\left\{\sup_{t\in[0,\tau_{n}\wedge T]}\left\|{\Phi(t)}\right\|^{2}+\int_{0}^{\tau_{n}\wedge T}\left|{\mathcal{A}\Phi(s)}\right|^{2}\mathrm{d}s\geq n\right\}\cap\left\{\sigma_{M}>T\right\}\right]$
$\leq\mathbb{P}\left[\sup_{t\in[0,\tau_{n}\wedge\sigma_{M}]}\left\|{\Phi(t)}\right\|^{2}+\int_{0}^{\tau_{n}\wedge\sigma_{M}}\left|{\mathcal{A}\Phi(s)}\right|^{2}\mathrm{d}s\geq n\right].$ (5.13)
We need to estimate this last probability. Let $T,M$, and $n$ be fixed. Let $\tau_{a},\tau_{b}$ be a pair of stopping times such that $0\leq\tau_{a}\leq\tau_{b}\leq\tau_{n}\wedge\sigma_{M}$. Applying Itô’s lemma to $\left\|{\Phi}\right\|^{2}$ and following the same argument as in (4), we have
$\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi(t)}\right\|^{2}\right]+2\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi(s)}\right|^{2}\mathrm{d}s\right]$
$\leq\mathbb{E}\left[\left\|{\Phi(\tau_{a})}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\left\langle\mathcal{B}(\Phi),\mathcal{A}\Phi\right\rangle}\right|\mathrm{d}s\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\left\langle\mathcal{R}(\Phi),\mathcal{A}\Phi\right\rangle}\right|\mathrm{d}s\right]$
$\quad+\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\sum_{k=1}^{\infty}\left\|{g_{k}(\Phi)}\right\|^{2}\mathrm{d}s\right]+\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left|{\sum_{k=1}^{\infty}\int_{\tau_{a}}^{t}2\left\langle g_{k}(\Phi),\Phi\right\rangle\mathrm{d}\beta_{k}}\right|\right].$
$=:\mathbb{E}\left[\left\|{\Phi(\tau_{a})}\right\|^{2}\right]+J_{1}+J_{2}+J_{3}+J_{4}.$ (5.14)
For the term $J_{1}$, we use (2.3) for $d=2$ and Young’s inequality to obtain
$J_{1}$ $\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\Phi}\right|^{\frac{1}{2}}\left\|{\Phi}\right\|\left|{\mathcal{A}\Phi}\right|^{\frac{3}{2}}\mathrm{d}s\right]\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\Phi}\right|^{2}\left\|{\Phi}\right\|^{4}\mathrm{d}s\right]+\frac{1}{2}\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi}\right|^{2}\mathrm{d}s\right].$
Next, we apply (2.8) for $d=2$ and Young’s inequality to estimate $J_{2}$ as
$J_{2}$ $\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left(1+\left|{\Phi}\right|^{\frac{3}{2}}\left|{\mathcal{A}\Phi}\right|^{\frac{1}{2}}\right)\left\|{\Phi}\right\|\left|{\mathcal{A}\Phi}\right|\mathrm{d}s\right]$
$\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi}\right\|^{2}+\left|{\Phi}\right|^{6}\left\|{\Phi}\right\|^{4}\mathrm{d}s\right]+\frac{1}{2}\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi}\right|^{2}\mathrm{d}s\right].$
By the assumption on $g$, we also have
$J_{3}\leq C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left(1+\left\|{\Phi}\right\|^{2}\right)\mathrm{d}s\right].$
For the last term, by the Burkholder–Davis–Gundy inequality, we have
$J_{4}\leq\frac{1}{2}\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi(t)}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left(1+\left\|{\Phi}\right\|^{2}\right)\mathrm{d}s\right].$
Substituting these estimates into (5.2) and rearranging the terms, we obtain
$\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi(t)}\right\|^{2}\right]+\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi(s)}\right|^{2}\mathrm{d}s\right]$
$\leq C\mathbb{E}\left[\left\|{\Phi(\tau_{a})}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left(1+\left|{\Phi}\right|^{6}\left\|{\Phi}\right\|^{2}\right)\left\|{\Phi}\right\|^{2}\mathrm{d}s\right].$ (5.15)
Note that by the definition of $\sigma_{M}$ in (5.10), we have
$\int_{0}^{\sigma_{M}}\left|{\Phi}\right|^{6}\left\|{\Phi}\right\|^{2}\mathrm{d}s\leq M,\;\text{ a.s.}$
We apply the Gronwall lemma for stochastic processes (Lemma A.4) to (5.2) to obtain
$\mathbb{E}\left[\sup_{t\in[0,\tau_{n}\wedge\sigma_{M}]}\left\|{\Phi(t)}\right\|^{2}+\int_{0}^{\tau_{n}\wedge\sigma_{M}}\left|{\mathcal{A}\Phi(s)}\right|^{2}\mathrm{d}s\right]\leq C_{T,M}\left(1+\left\|{\Phi_{0}}\right\|^{2}\right),$
where $C_{T,M}$ depends on $T$ and $M$, but is independent of $N$. Now, continuing from (5.2), by the Chebyshev inequality we have
$I_{2}$ $\leq\frac{1}{n}\mathbb{E}\left[\sup_{t\in[0,\tau_{n}\wedge\sigma_{M}]}\left\|{\Phi(t)}\right\|^{2}+\int_{0}^{\tau_{n}\wedge\sigma_{M}}\left|{\mathcal{A}\Phi(s)}\right|^{2}\mathrm{d}s\right]\leq\frac{C_{T,M}}{n}\left(1+\left\|{\Phi_{0}}\right\|^{2}\right).$ (5.16)
Altogether, from (5.11), (5.12), and (5.16), we deduce that for any $T\geq 1$,
$\mathbb{P}\left[\tau_{n}\leq T\right]\leq\frac{1}{M}\mathbb{E}\left[\int_{0}^{T\wedge\xi}\left|{\Phi(s)}\right|^{6}\left\|{\Phi(s)}\right\|^{2}\mathrm{d}s\right]+\frac{C_{T,M}}{n}\left(1+\left\|{\Phi_{0}}\right\|^{2}\right).$
First, we let $n\to\infty$ to obtain
$\lim_{n\to\infty}\mathbb{P}\left[\tau_{n}\leq T\right]\leq\frac{1}{M}\mathbb{E}\left[\int_{0}^{T\wedge\xi}\left|{\Phi(s)}\right|^{6}\left\|{\Phi(s)}\right\|^{2}\mathrm{d}s\right].$
Since this is true for any $M>0$, we conclude that $\lim_{n\to\infty}\mathbb{P}\left[\tau_{n}\leq T\right]=0$, thus $\mathbb{P}\left[\xi<\infty\right]=0$, as required in (5.9). ∎
Appendix A Auxiliary results
For convenience, we state several auxiliary results that are used in the paper. The first is a compactness theorem for stochastic processes satisfying a certain Cauchy criterion [24, 36]. The second is a generalised Poincaré-type inequalities, while the third is a general lemma concerning weak convergence in Banach spaces. The last result is a Gronwall-type lemma for stochastic processes [24].
Theorem A.1 ([24, 36]).
Suppose that $(X,\left|{\cdot}\right|_{X})$ and $(Y,\left|{\cdot}\right|_{Y})$ are Banach spaces with continuous embedding $Y\subset X$. Define the space
$\mathcal{E}(T):=C([0,T];X)\cap L^{2}(0,T;Y)$
with the norm
$\left|{v}\right|_{\mathcal{E}(T)}:=\left(\sup_{t\in[0,T]}\left|{v(t)}\right|_{X}^{2}+\int_{0}^{T}\left|{v(t)}\right|_{Y}^{2}\mathrm{d}t\right)^{\frac{1}{2}}.$
Let $\{v_{n}\}_{n\in\mathbb{N}}$ be a sequence of $Y$-valued stochastic processes so that for every $T>0$, $v_{n}\in\mathcal{E}(T)$ a.s. For any $M>1$ and $T>0$, define a collection of stopping times
$\mathcal{T}_{n}^{M,T}:=\left\{\tau\leq T:\left|{v_{n}}\right|_{\mathcal{E}(\tau)}\leq M+\left|{v_{n}(0)}\right|_{X}\right\},$
and let $\mathcal{T}_{n,m}^{M,T}:=\mathcal{T}_{n}^{M,T}\cap\mathcal{T}_{m}^{M,T}$.
  1. (i)
    Suppose that, for some $M>1$ and $T>0$, we have
    $\lim_{n\to\infty}\sup_{m\geq n}\,\sup_{\tau\in\mathcal{T}_{m,n}^{M,T}}\mathbb{E}\left[\left|{v_{n}-v_{m}}\right|_{\mathcal{E}(\tau)}\right]$ $=0,$ (A.1)
    $\lim_{S\to 0}\sup_{n\in\mathbb{N}}\,\sup_{\tau\in\mathcal{T}_{n}^{M,T}}\mathbb{P}\left[\left|{v_{n}}\right|_{\mathcal{E}(\tau\wedge S)}^{2}>\left|{v_{n}(0)}\right|_{X}^{2}+(M-1)^{2}\right]$ $=0.$ (A.2)
    Then there exists a stopping time $\tau$ with $\mathbb{P}(0<\tau\leq T)=1$, a process $v$ with $v(\,\cdot\,\wedge\tau)\in\mathcal{E}(\tau)$, and a subsequence $\{v_{n_{l}}\}$ such that
    $\left|{v_{n_{l}}-v}\right|_{\mathcal{E}(\tau)}\to 0,\;\text{ a.s. when $l\to\infty$,}$
    $\left|{v}\right|_{\mathcal{E}(\tau)}\leq M+\sup_{n\in\mathbb{N}}\left|{v_{n}(0)}\right|_{X}\;\text{ a.s.}.$
  2. (ii)
    If, in addition to (A.1) and (A.2), we also have
    $\sup_{n\in\mathbb{N}}\mathbb{E}\left[\left|{v_{n}(0)}\right|_{X}^{p}\right]<\infty\;\text{ for some $p\in[1,\infty)$},$ (A.3)
    then
    $\mathbb{E}\left[\left|{v}\right|_{\mathcal{E}(\tau)}^{p}\right]$ $\leq C_{p}\left(M^{p}+\sup_{n}\mathbb{E}\left[\left|{v_{n}(0)}\right|_{X}^{p}\right]\right).$ (A.4)
    Furthermore, there exists a sequence of measurable sets $\Omega_{l}\uparrow\Omega$ such that
    $\sup_{l\in\mathbb{N}}\mathbb{E}\left[\mathds{1}_{\Omega_{l}}\left|{v_{n_{l}}}\right|_{\mathcal{E}(\tau)}^{p}\right]$ $<\infty,$ (A.5)
    $\lim_{l\to\infty}\mathbb{E}\left[\mathds{1}_{\Omega_{l}}\left|{v_{n_{l}}-v}\right|_{\mathcal{E}(\tau)}^{p}\right]$ $=0.$ (A.6)
Lemma A.2 ([24]).
Suppose that $X$ is a separable Banach space and let $Y\subset X$ be a dense subset. Let $X^{\prime}$ be the dual of $X$ and denote the dual pairing between $u\in X$ and $v\in X^{\prime}$ by $\left\langle u,v\right\rangle$. Assume that $(E,\mathcal{E},\mu)$ is a finite measure space. Let $p\in(1,\infty)$. Suppose that $\{v_{n}\}_{n\in\mathbb{N}}$ is a bounded sequence in $L^{p}(E,X^{\prime})$ and
$\left\langle v_{n},y\right\rangle\to\left\langle v,y\right\rangle,\;\text{ $\mu$-a.e. for all $y\in Y$.}$ (A.7)
Then $v_{n}\overset{\ast}{\rightharpoonup}v$ in $L^{p}(E,X^{\prime})$, i.e., (A.7) holds for all $y\in X.$
Lemma A.3 (Poincaré-type inequalities).
Let $\mathcal{H}$ and $\mathcal{A}$ be defined as in Subsection 2.2. Let $\mathcal{H}_{n}:=\text{span}\{e_{1},e_{2},\ldots,e_{n}\}$, where $\{e_{k}\}_{k\in\mathbb{N}}$ is an orthonormal basis for $\mathcal{H}$ consisting of eigenfunctions of $\mathcal{A}$. Let $P_{n}:\mathcal{H}\to\mathcal{H}_{n}$ be the projection operator and let $Q_{n}=I-P_{n}$, where $I$ is the identity operator. Then for any real numbers $\alpha_{1}$ and $\alpha_{2}$, $\alpha_{1}<\alpha_{2}$, if $v\in\mathrm{D}(\mathcal{A}^{\alpha_{2}})$, then
$\left|{Q_{n}v}\right|_{\alpha_{1}}$ $\leq\lambda_{n}^{\alpha_{1}-\alpha_{2}}\left|{Q_{n}v}\right|_{\alpha_{2}},$ (A.8)
$\left|{P_{n}v}\right|_{\alpha_{2}}$ $\leq\lambda_{n}^{\alpha_{2}-\alpha_{1}}\left|{P_{n}v}\right|_{\alpha_{1}}.$ (A.9)
Lemma A.4 (Gronwall-type lemma for stochastic processes [24]).
Let $T>0$ be fixed. Suppose that $b,u,v,w:[0,T)\times\Omega\to\mathbb{R}$ are real-valued, non-negative stochastic processes. Let $\tau<T$ be a stopping time such that
$\mathbb{E}\left[\int_{0}^{\tau}(b(t)u(t)+w(t))\,\mathrm{d}t\right]<\infty.$
Moreover, assume that for a fixed constant $\kappa$,
$\int_{0}^{\tau}b(t)\,\mathrm{d}t\leq\kappa,\;\text{ a.s.}$
Suppose that there exists a constant $C_{0}$ such that for all stopping times $\tau_{a},\tau_{b}$ with $0\leq\tau_{a}<\tau_{b}\leq\tau$, we have
$\mathbb{E}\left[\sup_{s\in[\tau_{a},\tau_{b}]}u(s)+\int_{\tau_{a}}^{\tau_{b}}v(t)\,\mathrm{d}t\right]\leq C_{0}\mathbb{E}\left[u(\tau_{a})+\int_{\tau_{a}}^{\tau_{b}}(b(t)u(t)+w(t))\,\mathrm{d}t\right].$
Then
$\mathbb{E}\left[\sup_{s\in[0,\tau]}u(s)+\int_{0}^{\tau}v(t)\,\mathrm{d}t\right]\leq C\mathbb{E}\left[u(0)+\int_{0}^{\tau}w(t)\,\mathrm{d}t\right],$
where $C:=C(C_{0},T,\kappa)$.
Statements and declarations
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Funding statement
The authors acknowledge financial support through the Australian Research Council’s Discovery Projects funding scheme (projects DP200101866 and DP220101811).
Agus L. Soenjaya is supported by the Australian Government Research Training Program (RTP) Scholarship awarded at the University of New South Wales, Sydney.
Data availability
No data are associated with this article.